### 12 Mar 2026 v1

# Spectral finiteness, quantum norm continuity and classical points

Abstract
We prove various notions of uniform continuity for compact-quantum-group representations on Hilbert or Banach spaces equivalent to having finite spectrum, i.e. finitely many isotypic components. This generalizes the classical analogue for compact-group representations on Banach spaces, and relies in part on Riemann-Lebesgue-type decay properties for Fourier coefficients of elements in minimal tensor products with compact-quantum-group function algebras.
Key words: Riemann-Lebesgue; compact quantum group; conditional expectation; convolution product; injective tensor product; irreducible representation; norm-continuous representation; spectrum
MSC 2020: 46L67; 22D10; 16T05; 16T15; 47B01; 43A32; 46L85; 47L65
Introduction
The aim in the present note is to compare and relate two aspects of compact-quantum group unitary representations:
  • on the one hand analytic: continuity/decay properties mimicking, noncommutative-geometrically, classical norm-continuity for a unitary representation ${\mathbb{G}}\to{\mathbb{U}}(H)$ of a compact group $\mathbb{G}$} on a Hilbert space $H$;
  • on the other, representation-theoretic: the representation’s having finite spectrum, i.e. finitely many isotypic components.
Classically, the two properties (norm continuity and spectral finiteness) are indeed equivalent: [15, p.257] addresses this for connected second-countable groups, [27, Corollary 2] in the broader setup of arbitrary compact groups represented norm-continuously on Banach spaces, and [8, Theorem 3.10] provides a number of equivalent characterizations.
The compact quantum groups $\mathbb{G}$} to which the above uniformity/spectral finiteness discussion is to be extended are here as in [23, Definition 1.1.1]: unital $C^{*}$-algebras ${\mathcal{C}}({\mathbb{G}})$ (thought of as continuous-function algebras on the otherwise-nonexistent $\mathbb{G}$}) equipped with a coassociative unital $C^{*}$ morphism
${\mathcal{C}}({\mathbb{G}})\xrightarrow{\quad\Delta\quad}{\mathcal{C}}({\mathbb{G}})\underline{\otimes}{\mathcal{C}}({\mathbb{G}})\quad\left(\text{spatial $C^{*}$ tensor product}\right)$
with (the products’ linear spans) $\left({\mathcal{C}}({\mathbb{G}})\otimes 1\right)\Delta{\mathcal{C}}({\mathbb{G}})$ and $\left(1\otimes{\mathcal{C}}({\mathbb{G}})\right)\Delta{\mathcal{C}}({\mathbb{G}})$ both dense in ${\mathcal{C}}({\mathbb{G}})^{\underline{\otimes}2}$.
We recall briefly in Section˜1 below how unitary $\mathbb{G}$}-representations on Hilbert spaces $H$ are typically cast as unitary elements $U\in M({\mathcal{C}}({\mathbb{G}})\underline{\otimes}{\mathcal{K}}(H))$: multiplier algebra of the displayed minimal $C^{*}$ tensor product between the quantum function algebra ${\mathcal{C}}({\mathbb{G}})$ and the algebra of compact operators on $H$. That taken for granted, the quantization of the (uniformity=spectral finiteness) principle reads:
Theorem 0.1.
The following conditions on a unitary $\mathbb{G}$}-representation $U\in M({\mathcal{C}}({\mathbb{G}})\underline{\otimes}{\mathcal{K}}(H))$ on a Hilbert space $H$ are equivalent.
  1. (a)
    $U$ has finite spectrum.
  2. (b)
    The continuous adjoint action
    ${\mathcal{K}}(H)\xrightarrow{\quad U^{*}(1\otimes-)U\quad}{\mathcal{C}}({\mathbb{G}})\underline{\otimes}{\mathcal{K}}(H)$
    extends to a continuous action on ${\mathcal{L}}(H)$.
  3. (c)
    $U$ is norm-continuous, in the sense that $U\in{\mathcal{C}}({\mathbb{G}})\underline{\otimes}{\mathcal{L}}(H)$.
There are alternative possible notions of norm continuity for compact-quantum-group representations, beyond those featuring in Theorem˜0.1. [8, Definition 3.6(3)], for instance, expands the scope of the discussion to representations on Banach (as opposed to Hilbert) spaces $E$ in the sense of [8, Definition 3.1]: coassociative maps $E\xrightarrow{\rho}{\mathcal{C}}({\mathbb{G}})\otimes_{\lambda}E$ into the injective [12, §1.1] Banach-space tensor product satisfying the customary density condition (see ˜1-5 below). In that Banach-space context, said norm (or uniform) continuity is defined via [8, Lemma 3.7] as
(0-1) ${\mathcal{C}}({\mathbb{G}})^{*}\ni\psi\xmapsto[\quad\text{weak${}^{*}$-to-norm continuous}\quad]{\quad}\left(v\xmapsto{\quad}(\psi\otimes\operatorname{id})\rho v\right)\in{\mathcal{L}}(E).$
Spectral machinery extends to Banach-space representations, and one direct quantum counterpart to the classical coincidence of uniformity with spectrum finiteness is
Theorem 0.2.
For a compact quantum group $\mathbb{G}$} represented via $E\xrightarrow{\rho}{\mathcal{C}}({\mathbb{G}})\otimes_{\lambda}E$ on a Banach space $E$ the following are equivalent.
  1. (a)
    $\operatorname{\mathrm{spec}}\rho$ is finite.
  2. (b)
    $\rho$ is norm-continuous in the sense that ˜0-1 holds.
This reverses the implication recorded as [8, Lemma 3.8], but note that the hypothesis is arguably somewhat excessive: classically, one need not require the full power of ˜0-1, for in that case norm uniformity translates directly to the map in question being continuous when restricted to the space of pure states of ${\mathcal{C}}({\mathbb{G}})$ (i.e. $\mathbb{G}$} itself, with points identified with their respective Dirac-delta measures). This latter remark too has a quantum analogue, in the form of Theorem˜0.3 below. To make sense of the statement:
  • we reference the matrix coefficients $u^{\alpha}_{ij}\in{\mathcal{C}}({\mathbb{G}})$ (see Section˜1’s brief recollection) with $\alpha$ ranging over irreducible $\mathbb{G}$}-representations $\operatorname{Irr}({\mathbb{G}})$;
  • and will refer to a unitary $\mathbb{G}$}-representation $U$ as uniform≤1 (Definition˜1.3) provided ˜0-1’s map is assumed continuous only on the unit ball.
Theorem 0.3.
Let $U\in M({\mathcal{C}}({\mathbb{G}})\underline{\otimes}{\mathcal{K}}(H))$ be a unitary representation of a compact quantum group $\mathbb{G}$}.
  1. (1)
    If the Pontryagin dual [29, Theorem 3.3.2] $\Gamma:=\widehat{{\mathbb{G}}}$ has tempered decay in the sense that
    $\begin{gathered}\exists\left(C>0\right)\forall\left(\alpha\in\operatorname{Irr}({\mathbb{G}})\right)\left(\left\|u^{\alpha}\right\|_{\infty}^{{\mathcal{C}}({\mathbb{G}})}>C\right)\\ \left\|u^{\alpha}\right\|_{\infty}^{{\mathcal{C}}({\mathbb{G}})}:=\sup_{\begin{subarray}{c}v,w\in{\mathbb{C}}^{\dim\alpha}\\ \|v\|,\|w\|\leq 1\end{subarray}}\left\|\Braket{v\mid u^{\alpha}\mid w}:=\sum_{i,j=1}^{\dim\alpha}u^{\alpha}_{ij}\overline{v_{i}}w_{j}\right\|.\end{gathered}$
    $U$ is uniform≤1 if and only if it has finite spectrum.
  2. (2)
    In particular, said equivalence holds if $\mathbb{G}$} has a classical point, in the sense that ${\mathcal{C}}({\mathbb{G}})$ has a multiplicative state.
Counterexamples (see Theorem˜1.9) show that one cannot completely dispense with some type of decay control on the matrix coefficients of $\mathbb{G}$}.
Acknowledgments
Initial drafts have benefited from valuable input from M. Brannan.
1 Quantum uniform continuity in several guises, and consequences
Some of the usual operator-algebraic notation and conventions will be operative: $A\bullet B$ with $\bullet\in\{\otimes,\ \underline{\otimes},\ \overline$$\otimes$}\} stand for the ordinary algebraic, and spatial (or minimal or injective) [28, Definitions IV.4.8 and IV.5.1] $C^{*}$ and $W^{*}$ tensor products respectively (between operator algebras of context-appropriate type), M()M(-) is the multiplier algebra [2, Theorem II.7.3.1], ${\mathcal{L}}(-,-)$ and ${\mathcal{K}}(-,-)$ denote spaces of bounded and compact operators respectively between Hilbert/Banach spaces, and we refer freely to slice maps ([2, §§II.9.7.1, III.2.2.6], [19, Lemma 1.5], [30, §1], etc.)
$\begin{aligned} M\left(A\underline{\otimes}B\right)&\xrightarrow{\quad\psi\otimes\operatorname{id}\quad}M(B)\\ A\overline{\otimes}B&\xrightarrow{\quad\psi\otimes\operatorname{id}\quad}B\end{aligned},\quad\psi\in A^{*}\left(:=\text{continuous dual of $A$}\right)$
For compact quantum $\mathbb{G}$} with underlying function algebra ${\mathcal{C}}({\mathbb{G}})$ we work throughout with Woronowicz’s matrix coefficients ([34, §§1 and 4], [25, §1, pp.2-3], [23, §§1.3 and 1.4], etc.)
$u^{\alpha}_{ij}\in{\mathcal{O}}({\mathbb{G}}),\quad\begin{aligned} \alpha\in\operatorname{Irr}({\mathbb{G}})&:=\left\{\text{iso classes of simple ${\mathcal{O}}({\mathbb{G}})$-comodules}\right\}\\ 1\leq i,j\leq d_{\alpha}&:=\dim\alpha\end{aligned}$
spanning the (unique dense [18, Theorem 3.1.7]) Hopf *-algebra of representative functions on a compact quantum group $\mathbb{G}$}. Unitary $\mathbb{G}$}-representations on Hilbert spaces $H$ are as in [11, §2.2]:
$\text{unitary }U\in M({\mathcal{C}}({\mathbb{G}})\underline{\otimes}{\mathcal{K}}(H)),\quad(\Delta\otimes\operatorname{id})U=U_{13}U_{23},$
with $U_{ij}$ denoting $U$ applied on the $i^{th}$ and $j^{th}$ tensorands of the argument. Just as classically [13, Theorem 4.22], there is a rich Peter-Weyl theory of unitary-representation decompositions into irreducible summands isomorphic to the various $\alpha\in\operatorname{Irr}({\mathbb{G}})$ and corresponding spectral projections
$\left(\psi^{\alpha}\otimes\operatorname{id}\right)U^{*}=:P^{\alpha}=\left(P^{\alpha}\right)^{*}=\left(P^{\alpha}\right)^{2}\in{\mathcal{L}}(H)$
with $\psi^{\alpha}\in{\mathcal{C}}({\mathbb{G}})^{*}$ the functionals of [25, (4)], agreeing with the counit $\varepsilon$ of ${\mathcal{O}}({\mathbb{G}})$ on the span of $u^{\alpha}_{ij}$ and annihilating all $u_{k\ell}^{\alpha^{\prime}\neq\alpha}$. The range $\operatorname{\mathrm{im}}P^{\alpha}$ is the $\alpha$-isotypic component of $U$ (the largest subrepresentation isomorphic to a sum of copies of $\alpha$), and
$\operatorname{\mathrm{spec}}U:=\left\{\alpha\in\operatorname{Irr}({\mathbb{G}})\ :\ P^{\alpha}\neq 0\right\}$
is the spectrum of $U$.
It will be convenient to also assume the canonical positive operators
[…]
diagonal. One effect this will have is to render the $u^{\alpha}_{ij}$, (via [34, (5.14) and (5.15)] or [23, Theorem 1.4.3]), mutually orthogonal with respect to both versions of the inner product induced by the Haar state $h:=h_{{\mathbb{G}}}$ of [34, §4] or [23, §1.2]:
(1-1) $\Braket{x\mid y}_{r}:=h(x^{*}y)\quad\text{and}\quad\Braket{x\mid y}_{\ell}:=h(xy^{*}),$
the index indicating the position (left/right) of the product’s linear argument.
For a unital $C^{*}$-algebra $A$ an element $x\in{\mathcal{C}}({\mathbb{G}})\underline{\otimes}A$ has a Fourier-series decomposition
$x\sim\sum_{\begin{subarray}{c}\alpha\in\operatorname{Irr}({\mathbb{G}})\\ 1\leq i,j\leq d_{\alpha}\end{subarray}}u^{\alpha}_{ij}\otimes x^{\alpha}_{ij},$
the $\sim$” indicating heuristic rather than an equality in any formal sense (convergence issues being ignored); the rigorous version is rather
(1-2) $x^{\alpha}_{ij}:=\frac{q_{\alpha}}{F^{-1}_{\alpha ii}}(h\otimes\operatorname{id})\left(u^{\alpha*}_{ij}x\right)=\frac{q_{\alpha}}{F_{\alpha jj}}\left(h\otimes\operatorname{id})(xu^{\alpha*}_{ij}\right),$
the $u^{\alpha}_{ij}$ being assumed $h$-orthogonal as previously announced. This suggests the equally informal
$xx^{*}$ $\sim\sum_{\begin{subarray}{c}\alpha,i,j\\ \beta,k,l\end{subarray}}u^{\alpha}_{ij}u^{\beta*}_{kl}x^{\alpha}_{ij}x^{\beta*}_{kl}\xRightarrow{\quad}$
(1-3) $E\left(xx^{*}\right):=\left(h\otimes\operatorname{id}\right)(xx^{*})$ $=\sum_{\begin{subarray}{c}\alpha,i,j\\ \beta,k,l\end{subarray}}h\left(u^{\alpha}_{ij}u^{\beta*}_{kl}\right)x^{\alpha}_{ij}x^{\beta*}_{kl}$
[…]
The extent to which this last equality holds formally, with the right-hand side norm-convergent, is the focus of amply studied Fourier decay-at-infinity phenomena: quantum analogues of the familiar [4, §I.2] Riemann-Lebesgue lemma in classical Fourier analysis, as alluded to briefly in [1, §1, p.34], say. We revisit the matter in the course of the proof of Theorem˜0.1.
Note, first, an immediate consequence of Theorem˜0.1:
Corollary 1.1.
For any compact quantum group $\mathbb{G}$} and unital $C^{*}$-algebra $A$ every $\mathbb{G}$}-representation $U\in{\mathcal{C}}({\mathbb{G}})\underline{\otimes}A$ in $A$ in the sense of [17, Proposition 5.2] has finite spectrum.
Proof of Theorem˜0.1.
The equivalence ˜a $\Leftrightarrow$ ˜b is covered by [8, Theorem 3.16], while ˜a is plainly stronger than the other conditions: finite spectrum in fact entails
$U\in{\mathcal{O}}({\mathbb{G}})\otimes{\mathcal{L}}(H)\quad\left(\text{algebraic tensor product}\right).$
It remains to verify ˜c $\Rightarrow$ ˜a, which task the proof henceforth turns to. Reprising the notation of ˜1-3 for the usual conditional expectation [2, Definition II.6.10.1] ${\mathcal{C}}({\mathbb{G}})\underline{\otimes}{\mathcal{L}}(H)\xrightarrow[]{E}\mathrel{\mkern-14.0mu}\rightarrow{\mathcal{L}}(H)$, Proposition˜1.2 below ensures that that equation in fact holds as written, with a norm-convergent right-hand side. Applied to $x:=U$, ˜1-3 reads
$1=E(1)=E\left(xx^{*}\right)$ $=\sum_{\alpha,j}\frac{1}{q_{\alpha}}F_{\alpha jj}\sum_{i}x^{\alpha}_{ij}x^{\alpha*}_{ij}$
$=\sum_{\alpha}\left(\sum_{j}\frac{1}{q_{\alpha}}F_{\alpha jj}\right)P^{\alpha}$
$=\sum_{\alpha}P^{\alpha}$
as norm convergence, where
(1-4) $P^{\alpha}:=(\psi^{\alpha}\otimes\operatorname{id})U^{*}\in{\mathcal{L}}(H),\quad{\mathcal{C}}({\mathbb{G}})\ni u^{\beta}_{ij}\xmapsto{\quad\psi^{\alpha}\quad}\delta_{\alpha\beta}\delta_{ii}\in{\mathbb{C}}$
are the respective projections (introduced in [25, Theorem 1.5] as $E^{\alpha}$) onto the $\alpha$-isotypic components (cf. [3, remarks preceding Proposition 13]) of the representation $U$. Said norm convergence of mutually orthogonal projections, naturally, requires that only finitely many be non-zero. $\blacksquare$
The proof of the auxiliary Proposition˜1.2 below will precisely parallel its analogue [26, Lemma 5.2] in the context of reduced crossed products [6, Definition 4.1.4] resulting from actions by countable (classical, discrete) groups on $C^{*}$-algebras (so in our language, that result’s $\mathbb{G}$} would be an abelian, i.e. dual-classical, compact quantum group). That source leverages weaker convergence into its stronger, norm counterpart via a device credited there to G. Elliott and U. Haagerup and reliant on Dini’s [14, Theorem 12.1] [^1] Said invocation of Dini, typically stated for sequences of real-valued functions on compact Hausdorff spaces, appears to be the one aspect of the proof of [26, Lemma 5.2] relying in any substantive fashion on any countability hypotheses. We will see in the course of the proof of Proposition 1.2 that such countability assumptions can be dispensed with. (see also [20, Lemma 3.1] for essentially the same gadget).
Proposition 1.2.
For a compact quantum group $\mathbb{G}$}, unital $C^{*}$-algebra $A$ and $x\in{\mathcal{C}}({\mathbb{G}})\underline{\otimes}A$ we have the norm-convergent expansions
$E(xx^{*})=\sum_{\alpha,i,j}\frac{1}{q_{\alpha}}F_{\alpha jj}x^{\alpha}_{ij}x^{\alpha*}_{ij}\quad\text{and}\quad E(x^{*}x)=\sum_{\alpha,k,\ell}\frac{1}{q_{\alpha}}F_{\alpha kk}x^{\alpha*}_{k\ell}x^{\alpha}_{k\ell}$
for the respective values of the expectation ${\mathcal{C}}({\mathbb{G}})\underline{\otimes}A\xrightarrow{E:=h\otimes\operatorname{id}}A$.
Proof.
With nothing of any substance distinguishing the two branches, we address the first claim only. Write
  • $A\xrightarrow{\pi}{\mathcal{L}}(H)$ for a representation of $A$;
  • ${\mathcal{C}}({\mathbb{G}})\xrightarrow{\pi_{h}}{\mathcal{L}}\left(L^{2}({\mathbb{G}})\right)$ for the GNS representation of the Haar state $h\in{\mathcal{C}}({\mathbb{G}})^{*}$: the completion of ${\mathcal{O}}({\mathbb{G}})$ under the inner product $\Braket{-\mid-}_{r}$ of ˜1-1;
  • and
    $H\ni\xi\xmapsto[\quad\begin{subarray}{c}\alpha\in\operatorname{Irr}({\mathbb{G}})\\ 1\leq i,j\leq d_{\alpha}\end{subarray}\quad]{\quad T^{\alpha}_{ij}\quad}u^{\alpha*}_{ij}\otimes\xi\in L^{2}({\mathbb{G}})\otimes H.$
Observe, next, the convergence
$\sum_{\alpha,i,j}\frac{q_{\alpha}}{F_{\alpha jj}}T^{\alpha}_{ij}T^{\alpha*}_{ij}=1\in{\mathcal{L}}\left(L^{2}({\mathbb{G}})\otimes H\right)$
in the $\sigma$-strong topology of [28, Definition II.2.3] (the strongest of the locally convex sub-norm topologies examined in [28, §II.2]): the individual summands $\frac{q_{\alpha}}{F_{\alpha jj}}T^{\alpha}_{ij}T^{\alpha*}_{ij}$ are (by [34, (5.14)] again) mutually orthogonal projections whose ranges exhaust the underlying Hilbert space.
Note also:
$\forall\left(\alpha\in\operatorname{Irr}({\mathbb{G}})\right)\forall\left(1\leq i,j\leq d_{\alpha}\right)\bigg(\pi x^{\alpha}_{ij}=\frac{q_{\alpha}}{F_{\alpha jj}}T^{\mathbf{1}*}_{11}\left(\pi_{h}\otimes\pi\right)(x)T^{\alpha}_{ij}\bigg)$
($\mathbf{1}\in\operatorname{Irr}({\mathbb{G}})$ denoting the trivial representation). As in the aforementioned [26, Lemma 5.2], this now entails the pointwise convergence
$\pi E(xx^{*})=\sum_{\alpha,i,j}\frac{F_{\alpha jj}}{q_{\alpha}}\pi\left(x^{\alpha}_{ij}x^{\alpha*}_{ij}\right)$
on the space of states on $A$ inherited from the weak ones on ${\mathcal{L}}(H)$ via $\pi$. We can arrange for this to be the entire state space ${\mathcal{S}}(A)$ (take for $\pi$ the universal representation [24, §3.7.6]), hence also uniform convergence by the version of Dini’s theorem proven in [22]: valid for monotone nets of functions valued in arbitrary uniform spaces (monotonicity being defined appropriately to that context) rather than real-valued-function sequences. $\blacksquare$
Recall next the notion of $\mathbb{G}$}-representation on a Banach space $E$ for a compact quantum group $\mathbb{G}$}:
[…]
One again has spectral idempotents $P^{\alpha}:=(\psi^{\alpha}\otimes\operatorname{id})\circ\rho\in{\mathcal{L}}(E)$ analogous to ˜1-4 and, by [8, Proposition], the more familiar unitary representations $U\in M({\mathcal{C}}({\mathbb{G}})\underline{\otimes}{\mathcal{K}}(H))$ can be recast as representations on the underlying Banach spaces by
(1-6) $H\ni v\xmapsto{\quad}U^{*}(1\otimes v)\in{\mathcal{C}}({\mathbb{G}})\otimes_{\lambda}H.$
Proof of Theorem˜0.2.
While [8, Lemma 3.8] notes only that having finite spectrum implies ˜0-1, the converse is also straightforward: weak-to-norm continuity is an extremely strong condition to impose, holding only in trivial circumstances.
To elaborate, consider a linear map $E\xrightarrow{T}F$ between locally convex [16, §18.1] topological vector spaces. Its continuity with respect to
  • the weak (or simple) topology ($\mathfrak{T}_{s}$ in [16, §20.2]);
  • and the given topology on $F$, assuming some origin neighborhood $U$ contains no non-trivial linear subspaces of $F$ (so in particular for normable $F$)
automatically entails the finite dimensionality of $\operatorname{\mathrm{im}}T$. This is little more than an unpacking of the hypotheses: there must be finitely many continuous functionals $f_{i}\in E^{*}$ with
$\forall\left(v\in\bigcap_{i}\ker f_{i}\right)\left(Tv\in U\right).$
One can scale $v$ at will, so the assumption on $U$ forces $T$ to vanish on the finite-codimensional $\bigcap_{i}\ker f_{i}$. $\blacksquare$
In view of the preceding proof, rendering Theorem˜0.2 immediate by means of unrestrained weak continuity, one might attempt adjacent refinements to uniformity.
Definition 1.3.
A Banach-space representation ˜1-5 of a compact quantum group is boundedly uniform or uniform≤1 if the continuity in ˜0-1 holds on norm-bounded subsets of ${\mathcal{C}}({\mathbb{G}})^{*}$ (equivalently, on any one norm-bounded origin neighborhood, e.g. the unit ball). $\blacklozenge$
For classical $\mathbb{G}$} bounded and plain uniformity are equivalent: as noted in the Introduction following Theorem˜0.2, uniformity≤1 already implies spectrum finiteness and hence [8, Lemma 3.8] also ordinary uniformity. It is not inapposite, then, to ask whether the equivalence survives quantization.
As further motivation and surrounding context for Theorem˜0.3, recall [7, Remark 1.3] to the effect that the classical [27, Corollary 2] characterization of norm-continuous compact-group representations as precisely those with finite spectra is recovered by non-classical means from [7, Theorem 0.1] for compact quantum groups $\mathbb{G}$} with a classical point (i.e. with ${\mathcal{C}}({\mathbb{G}})$ admitting a multiplicative state); by extension, the same goes for Theorem˜0.1 (which removes the classical-point constraint). Theorem˜0.3 describes how, upon assuming a small amount of classical behavior, said classical result replicates in terms of what is perhaps the most straightforward analogue of quantum uniform continuity for a unitary $\mathbb{G}$}-representation.
An initial remark will help set the stage for further investigation.
Lemma 1.4.
A unitary representation $U\in M({\mathcal{C}}({\mathbb{G}})\underline{\otimes}{\mathcal{K}}(H))$ of a compact quantum group $\mathbb{G}$} is uniform≤1 as a Banach-space representation if and only if either one of
${\mathcal{C}}({\mathbb{G}})^{*}\ni\psi\xmapsto{\quad}(\psi\otimes\operatorname{id})U\text{ or }(\psi\otimes\operatorname{id})U^{*}\in{\mathcal{L}}(H)$
is weak-to-norm continuous on the unit ball of ${\mathcal{C}}({\mathbb{G}})^{*}$.
Proof.
For the latter map this is immediate given the conversion procedure ˜1-6 from unitary to Banach-space representations, while the passage between the two is effected by $\left(\left(\psi\otimes\operatorname{id}\right)U\right)^{*}=\left(\psi^{*}\otimes\operatorname{id}\right)U^{*}$ for $\psi^{*}:=\overline{\psi\left(\bullet^{*}\right)}$. $\blacksquare$
The device isolated by Proposition˜1.6 below, implicitly also driving the proof of [30, Proposition 2.2], will allow some symmetrization in the condition on $U$ imposed by Lemma˜1.4. Some vocabulary will help compress the statement.
Definition 1.5.
For von Neumann algebras $M$ and $N$ and a property $\mathcal{P}$} that a map $M^{*}\to N$ may or may not have, $x\in M\overline{\otimes}N$ is said to have property $\mathcal{P}$$_{\ell}$ or to be $\mathcal{P}$$_{\ell}$ (for “left”) if the slice map
$M^{*}\ni\varphi\xmapsto{\quad}\left(\varphi\otimes\operatorname{id}\right)x\in N$
induced by $x$ is $\mathcal{P}$}.
The right-handed mirror is phrased in the expected fashion: ${\mathcal{P}}_{r}$ in place of $\mathcal{P}$$_{\ell}$, for the package involving $\operatorname{id}\otimes\varphi$ in place of $\varphi\otimes\operatorname{id}$. The terminology further applies in its guessable $C^{*}$ incarnation: $x\in M\underline{\otimes}N$. $\blacklozenge$
One relevant property $\mathcal{P}$} is the uniformity of ˜0-1 and (of more interest here) its bounded variant of Definition˜1.3.
Proposition 1.6.
For $x\in M\overline{\otimes}N$ in the $W^{*}$ tensor product of two von Neumann algebras the following conditions are equivalent.
  1. (a)
    $x$ is uniform≤1,ℓ.
  2. (b)
    $x$ is uniform≤1,ℓ with respect to (bounded) nets of weak functionals converging to 0.
  3. (c)
    As in ˜b, for nets of weak functionals instead.
  4. (d)
    As in ˜a, for nets of functionals of the form
    $M\ni T\xmapsto{\quad\omega_{v,w}\quad}\Braket{v\mid Tw}\in{\mathbb{C}}$
    for $v,w\in H$ in a Hilbert space on which $M$ is represented faithfully as $M\leq{\mathcal{L}}(H)$.
  5. (e)
    All or any of the above, in the right-handed version.
Proof.
The implications
[…]
are all formal.
Note that ˜a concerns all of $M^{*}$, whereas the other $\bullet_{\ell}$ specialize to (at least) weak-continuous functionals. The proof of [30, Proposition 2.2], in present language, shows effectively that the strongest left-handed statement ˜a is implied by its right-handed counterpart phrased for nets $(\psi_{\lambda})_{\lambda}\subset N_{*,\leq 1}$ (unit ball of the predual of $N$) assumed convergent $N^{*}_{\leq 1}$. This is easily recast as ˜br though: the net in question is certainly Cauchy and hence $(\varphi_{\lambda}-\varphi_{\lambda^{\prime}})_{\lambda,\lambda^{\prime}}$ is 0-convergent; consequently,
[…]
˜c $\Rightarrow$ ˜b by weak-in-weak norm density [28, Theorem II.2.6(iii)].
In order to complete the proof of the proposition as a whole by confirming that ˜er implies ˜a, it remains to observe that the aforementioned net $(\psi_{\lambda})_{\lambda}\subset N_{*,\leq 1}$ can be chosen so as to range over any set ${\mathcal{S}}\subset N_{\leq 1}^{*}$ detecting norms in $N$ in the following sense:
$\forall\left(T\in N\right)\forall\left(\varepsilon>0\right)\exists\left(\varphi\in{\mathcal{S}}\right)\left(|\varphi T|\geq\|T\|-\varepsilon\right).$
Plainly, the set $\mathcal{S}$} of all functionals $\omega_{v,w}$ with norm-$(\leq 1)$ $v,w\in H$ will do. $\blacksquare$
The aforementioned symmetrization afforded by Proposition˜1.6 refers to switching Lemma˜1.4’s criterion to $U$’s right-hand leg. For $v\in H$ we write
(1-7) $\operatorname{\mathrm{supp}}v:=\left\{\alpha\in\operatorname{Irr}({\mathbb{G}})\ :\ P_{U}^{\alpha}v\neq 0\right\}$
for the support of $v$ with respect to the representation $U$, and $\operatorname{\mathrm{supp}}v\to\infty$ for varying $v$ means that the support eventually misses every finite subset of $\operatorname{Irr}({\mathbb{G}})$.
Corollary 1.7.
A unitary representation $U\in M({\mathcal{C}}({\mathbb{G}})\underline{\otimes}{\mathcal{K}}(H))$ of a compact quantum group $\mathbb{G}$} is uniform≤1 as a Banach-space representation if and only if
$\left(\operatorname{\mathrm{supp}}v\to\infty\wedge\operatorname{\mathrm{supp}}w\to\infty\right)\ \xRightarrow{\quad}\ (\operatorname{id}\otimes\omega_{v,w})U\xrightarrow[\quad\quad]{\quad\text{norm}\quad}0$
for $(v,w)\in\left(H_{\leq 1}\right)^{2}$ and $\omega_{v,w}:=\Braket{v\mid\bullet w}$.
Proof.
Lemma˜1.4 and Proposition˜1.6’s equivalence ˜a $\Leftrightarrow$ ˜dr equate uniformity≤1 with the weak-norm convergence of
$\omega_{v,w}\xmapsto{\quad}\left(\operatorname{id}\otimes\omega_{v,w}\right)U\in{\mathcal{L}}(H),\quad v,w\in H_{\leq 1}.$
We may thus assume weakly convergent nets
$\left(v_{\lambda}\right)_{\lambda}\xrightarrow[\quad\lambda\quad]{\quad}v\quad\text{and}\quad\left(w_{\lambda}\right)_{\lambda}\xrightarrow[\quad\lambda\quad]{\quad}w$
in the unit ball $H_{\leq 1}$, whence $\omega_{v_{\lambda},w_{\lambda}}\to\omega_{v,w}$. Approximate $v$ and $w$ arbitrarily well by finitely supported vectors $\overline{v}$ and $\overline{w}$ respectively, assuming also that
$\operatorname{\mathrm{supp}}\left(v_{\lambda}-\overline{v}\right)\to\infty\quad\text{and}\quad\operatorname{\mathrm{supp}}\left(w_{\lambda}-\overline{w}\right)\to\infty.$
Because finite truncations
$\sum_{\alpha\in{\mathcal{F}}}\left(1\otimes P^{\alpha}\right)U,\quad\text{finite }{\mathcal{F}}\subseteq\operatorname{Irr}({\mathbb{G}})$
belong to the algebraic tensor product ${\mathcal{O}}({\mathbb{G}})\otimes{\mathcal{L}}(H)$, the norm convergence of the truncated
$\left(\operatorname{id}\otimes\omega_{\bullet}\right)U,\quad\bullet\in\left\{(\overline{v},\overline{w}),\ \left(v_{\lambda}-\overline{v},\overline{w}\right),\ \left(\overline{v},w_{\lambda}-\overline{w}\right)\right\}$
is not at issue. This confirms the equivalence
$\left(\operatorname{id}\otimes\omega_{v_{\lambda},w_{\lambda}}\right)U\xrightarrow[\quad\lambda\quad]{\quad\text{norm}\quad}\left(\operatorname{id}\otimes\omega_{v,w}\right)U\quad\xLeftrightarrow{\quad}\quad\left(\operatorname{id}\otimes\omega_{v_{\lambda}-\overline{v},w_{\lambda}-\overline{w}}\right)U\xrightarrow[\quad\lambda\quad]{\quad\text{norm}\quad}0,$
finishing the proof. $\blacksquare$
Proof of Theorem˜0.3.
  1. ˜1 Assume infinitely many spectral projections $P^{\alpha}$ of $U$ non-zero, entailing the corresponding non-vanishing Fourier coefficients $x^{\alpha}_{ij}$, $1\leq i,j\leq\dim\alpha$ of $x:=U$ defined by ˜1-2 (having fixed the $u^{\alpha}$-engendering bases of the $\alpha$ as explained preceding ˜1-1, so as to ensure the $u^{\alpha}_{ij}$ are mutually orthogonal). The hypothesis ensures the existence of norm-1
    $v_{0}=\left(v_{0i}\right)_{1\leq i\leq\dim\alpha}\quad\text{and}\quad w_{0}=\left(w_{0i}\right)_{1\leq i\leq\dim\alpha}\quad:\quad\left\|\sum_{i,j=1}^{\dim\alpha}u^{\alpha}_{ij}\overline{v_{0i}}w_{0j}\right\|>C.$
    To procure suitable $\omega_{v,w}$ amenable to Corollary˜1.7, set
    $v:=\sum_{i=1}^{\dim\alpha}v_{0i}e_{i}\quad\text{and}\quad w:=\sum_{i=1}^{\dim\alpha}w_{0i}e_{i},\quad e_{i}\in\operatorname{\mathrm{im}}x^{\alpha}_{ii}\leq\operatorname{\mathrm{im}}P^{\alpha}\text{ orthonormal}.$
    We have
    $\left\|(\operatorname{id}\otimes\omega_{v,w})U\right\|=\left\|\sum_{i,j=1}^{\dim\alpha}u^{\alpha}_{ij}\overline{v_{0i}}w_{0j}\right\|>C,$
    affording the desired conclusion that uniformity≤1 must fail.
  2. ˜2 $\mathbb{G}$} has a classical point precisely when the counit $\varepsilon$ has a continuous extension to all of ${\mathcal{C}}({\mathbb{G}})$ (see [10, paragraph preceding Theorem 2.2]; also Lemma˜1.8 below for additional equivalent characterizations), whence a state assigning the value 11 to every $u^{\alpha}_{ii}$. Those elements, then, must all have norm 1.
$\blacksquare$
In reference to the property featuring in [7, Theorem 0.1] of compact quantum groups having at least one classical point (one version of the above-mentioned “partial classical behavior”), Lemma˜1.8 records a simple remark for whatever future-reference purpose it may serve. We will be referring to the usual convolution product [21, Definition 1.4.1], denoted here simply by $\cdot$” or juxtaposition, and defined as the precomposition
${\mathcal{O}}({\mathbb{G}})^{*\otimes 2}\xrightarrow{\quad\circ\Delta\quad}{\mathcal{O}}({\mathbb{G}})^{*}$
on the algebraic dual of a coalgebra. Note also that it (or rather its appropriate extension) makes the state space ${\mathcal{S}}({\mathcal{C}}({\mathbb{G}}))$ of the $C^{*}$-algebra ${\mathcal{C}}({\mathbb{G}})$ into a compact Hausdorff [9, Theorem V.3.1] semigroup under its weak topology. The same goes for the unit ball ${\mathcal{C}}({\mathbb{G}})^{*}_{\leq 1}$ in the continuous dual of ${\mathcal{C}}({\mathbb{G}})$: the norm dual to $\underline{\otimes}$ being cross [32, Proposition T.3.11],
$\|\varphi\|,\ \|\psi\|\leq 1\xRightarrow{\quad}\|\varphi\cdot\psi:=(\varphi\otimes\psi)\circ\Delta\|\leq 1.$
The compact semigroups
${\mathcal{S}}:={\mathcal{S}}_{{\mathbb{G}}}:={\mathcal{S}}({\mathcal{C}}({\mathbb{G}}))\quad\text{and}\quad{\mathcal{K}}:={\mathcal{K}}_{{\mathbb{G}}}:={\mathcal{C}}({\mathbb{G}})^{*}_{\leq 1}$
will feature again below.
Lemma 1.8.
For a compact quantum group with underlying function algebra ${\mathcal{C}}({\mathbb{G}})$ the following conditions are equivalent.
  1. (a)
    $\mathbb{G}$} has a classical point, in the sense that ${\mathcal{C}}({\mathbb{G}})$ has a multiplicative state.
  2. (b)
    The counit of the Hopf *-algebra ${\mathcal{O}}({\mathbb{G}})$ extends continuously to ${\mathcal{C}}({\mathbb{G}})$.
  3. (c)
    One of the semigroups ${\mathcal{S}}_{{\mathbb{G}}}$, ${\mathcal{K}}_{{\mathbb{G}}}$ or ${\mathcal{C}}({\mathbb{G}})^{*}$ has an invertible element.
  4. (d)
    One of the semigroups ${\mathcal{S}}_{{\mathbb{G}}}$ or ${\mathcal{K}}_{{\mathbb{G}}}$ has an element $\psi$ with either multiplication map $\psi\cdot$ or $\cdot\psi$ bijective.
  5. (e)
    One of the semigroups ${\mathcal{S}}_{{\mathbb{G}}}$ or ${\mathcal{K}}_{{\mathbb{G}}}$ has an element $\psi$ with either multiplication map $\psi\cdot$ or $\cdot\psi$ surjective.
Proof.
We have
[…]
and
[…]
  1. Claim: the unit of any of the semigroups $\mathcal{S}$}, $\mathcal{K}$} or ${\mathcal{C}}({\mathbb{G}})^{*}$, if it exists, must be the counit. That the variants of the claim are mutually equivalent follows from the fact that arbitrary continuous functionals are expressible as linear combinations of at most four states (e.g. by [28, Theorem III.4.2(ii)]). Having thus reduced the problem to its ${\mathcal{C}}({\mathbb{G}})^{*}$ variant, write $\psi$ for the assumed unit of that convolution monoid and
    $\varepsilon_{{\mathcal{F}}}:=\begin{cases}\varepsilon&\text{ on }\operatorname{\mathrm{span}}\left\{u^{\alpha}_{ij}\ :\ \alpha\in{\mathcal{F}}\right\}\\ 0&\text{ on }\operatorname{\mathrm{span}}\left\{u^{\beta}_{ij}\ :\ \beta\not\in{\mathcal{F}}\right\}\end{cases},\quad\text{finite }{\mathcal{F}}\subseteq\operatorname{Irr}({\mathbb{G}}).$
    The $\varepsilon_{{\mathcal{F}}}$ being restrictions to ${\mathcal{O}}({\mathbb{G}})$ of continuous functionals (i.e. members of ${\mathcal{C}}({\mathbb{G}})^{*}$) by [34, (5.37) and surrounding discussion],
    $\psi=\lim_{{\mathcal{F}}\nearrow}\psi\cdot\varepsilon_{{\mathcal{F}}}=\lim_{{\mathcal{F}}\nearrow}\varepsilon_{{\mathcal{F}}}=\varepsilon$
    in the weak topology on the full linear dual ${\mathcal{O}}({\mathbb{G}})^{*}$.
[…]
    $\psi^{{n_{\lambda,\lambda^{\prime}}}}\cdot\in\left(\textsc{Cont}\left({\mathcal{K}}\to{\mathcal{K}}\right),\ \text{uniform topology}\right)$
    for $\mathcal{K}$}: this follows from the identification [5, Proposition 7.1.5]
    $\textsc{Cont}\left({\mathcal{K}}\times{\mathcal{K}}\to{\mathcal{K}}\right)\cong\textsc{Cont}\left({\mathcal{K}}\to\textsc{Cont}\left({\mathcal{K}}\to{\mathcal{K}}\right)\right).$
    Consequently ($\sim$” meaning “close”):
[…]
    The net $\left(\psi^{\left(n_{\lambda^{\prime}}-n_{\lambda}\right)}\right)_{\lambda^{\prime}>\lambda}$ will thus have a subnet converging to the identity, which must be a counit for ${\mathcal{C}}({\mathbb{G}})$ by the claim preceding the current portion of the argument.
$\blacksquare$
Theorem˜0.3 leaves unaddressed the now-inevitable question of whether the decay condition can be removed entirely. Due to (purely quantum) peculiarities, it turns out it cannot. To frame the result formally as Theorem˜1.9 below, we need to recollect some language attendant to quantum-group growth/decay considerations.
  • A finitely generated (f.g.) discrete quantum group is the Pontryagin dual $\Gamma:=\widehat{{\mathbb{G}}}$ for compact quantum $\mathbb{G}$} whose underlying Hopf *-algebra ${\mathcal{O}}({\mathbb{G}})$ is f.g. as an algebra.
  • The word length function [31, Example 3.2(ii)] $\ell$ on $\Gamma$ above with respect to a finite generating set ${\mathcal{F}}\subseteq\operatorname{Irr}({\mathbb{G}})$ is
    $\operatorname{Irr}({\mathbb{G}})\ni\alpha\xmapsto{\quad\ell=\ell_{{\mathcal{F}}}\quad}\min\left\{k\in{\mathbb{Z}}_{\geq 0}\ :\ \exists\left(\alpha_{i}\in{\mathcal{F}},\ 1\leq i\leq k\right)\left(\alpha\leq\alpha_{1}\otimes\cdots\otimes\alpha_{k}\right)\right\}.$
    We extend the notation $\ell(-)$ to elements of ${\mathcal{O}}({\mathbb{G}})$: one such has length $k$ by definition if it is a linear combination of matrix coefficients $u^{\alpha}_{ij}$ with $\ell(\alpha)=k$.
  • $\Gamma$ has the property of rapid decay (RD) [31, Proposition and Definition 3.5] (with respect to $\ell_{{\mathcal{F}}}$) if
    $\left(\forall x\in{\mathcal{O}}({\mathbb{G}})\right)\left(\|x\|\leq P(\ell(x))\|x\|_{2}\right),\quad\|-\|_{2}:=h\left(-^{*}\cdot-\right)^{1/2}$
    for some polynomial $P$.
  • The terminology applies here (somewhat unconventionally) also “locally” or relatively, to families of representations: given any finite ${\mathcal{F}}\subseteq\operatorname{Irr}({\mathbb{G}})$ one might consider sequences $\alpha_{n}$ exhibiting rapid decay with respect to $\ell:=\ell_{{\mathcal{F}}}$ in a sense the reader will have no difficulties extrapolating from the global version just sketched.
The following result shows (e.g. via Example˜1.10) that there do exist uniform≤1 representations with infinite spectrum.
Theorem 1.9.
Let $\mathbb{G}$} be a compact quantum group, ${\mathcal{F}}\subseteq\operatorname{Irr}({\mathbb{G}})$ a finite family and $\left(\alpha_{n}\right)_{n}\subset\operatorname{Irr}({\mathbb{G}})$ with property RD with respect to $\ell_{{\mathcal{F}}}$ and trivial $F^{\alpha_{n}}$.
If $\dim\alpha_{n}$ increase exponentially while the lengths $\ell\left(\alpha_{n}\right)$ increase polynomially, the unitary representation $\bigoplus_{n}\alpha_{n}$ is uniform≤1.
Proof.
Write $H:=\bigoplus_{n}H^{\alpha_{n}}$ for the carrier Hilbert space of the direct-sum representation $U$ in question. We have to argue, per Corollary˜1.7, that for $v\in H$ with
(1-8) $\|v\|=1\ \wedge\ \operatorname{\mathrm{supp}}v\subseteq\left\{\alpha_{m}\ :\ m\gg 0\right\}$
($\operatorname{\mathrm{supp}}$ defined as in ˜1-7) the norm $\left\|(\operatorname{id}\otimes\omega_{v,v})U\right\|$ will be as small as desired provided only that $m\gg 0$ is indeed sufficiently large. The unimodularity constraint $F^{\alpha}=1$ ensures [34, (5.14), (5.15)] that the $u^{\alpha_{n}}_{ij}$ (are mutually orthogonal and) have $\|-\|_{2}$ norms $\frac{1}{\sqrt{\dim\alpha_{n}}}$, hence the norm estimate
˜1-8 $\xRightarrow{\quad}\ \exists\left(c>1\right)\left(\left\|(\operatorname{id}\otimes\omega_{v,v})U\right\|_{2}\leq\max_{m:\alpha_{m}\in\operatorname{\mathrm{supp}}v}\frac{1}{\sqrt{\dim\alpha_{m}}}\leq\frac{1}{c^{m_{0}}}\right)$
$\xRightarrow{\quad\text{RD}\quad}\exists\left(\text{polynomial }P\right)\left(\left\|(\operatorname{id}\otimes\omega_{v,v})U\right\|\leq\frac{P(m_{0})}{c^{m_{0}}}\right)$
$\left(\text{with }m_{0}:=\min\left\{m\ :\ \alpha_{m}\in\operatorname{\mathrm{supp}}v\right\}\right)$
confirming the desired conclusion. $\blacksquare$
Example 1.10.
Examples of compact quantum groups verifying the hypotheses of Theorem˜1.9 are well known and have been much studied (from various perspectives, RD-related or not): both
  • the free orthogonal group ${\mathbb{G}}:={\mathbb{O}}^{+}(n)$, $n\geq 3$ (underlying function algebra ${\mathcal{C}}({\mathbb{G}})=A_{o}(I_{n})$ in [31, §4.3]), with ${\mathcal{O}}({\mathbb{G}})$ freely generated by a unitary $n\times n$ matrix $u^{\alpha}=\left(u^{\alpha}_{ij}\right)$, $1\leq i,j\leq n$ of self-adjoint elements;
  • and the free unitary group ${\mathbb{G}}:={\mathbb{U}}^{+}(n)$, $n\geq 3$, with ${\mathcal{O}}({\mathbb{G}})$ freely generated by unitary $u^{\alpha}\in{\mathcal{O}}({\mathbb{G}})\otimes M_{n}$ such that $\overline{u}^{\alpha}:=\left(u^{\alpha*}_{ij}\right)_{i,j}$ is also unitary
will do by [31, Theorems 4.9 and 4.10] respectively. $\blacklozenge$

### 13 Mar 2026 v2

@@
 There are alternative possible notions of norm continuity for compact-quantum-group representations, beyond those featuring in Theorem˜0.1.
-[8, Definition 3.6(3)], for instance, expands the scope of the discussion to representations on Banach (as opposed to Hilbert) spaces $E$ in the sense of [8, Definition 3.1]: coassociative maps $E\xrightarrow{\rho}{\mathcal{C}}({\mathbb{G}})\otimes_{\lambda}E$ into the injective [12, §1.1] Banach-space tensor product satisfying the customary density condition (see ˜1-5 below).
+[8, Definition 3.6(3)], for instance, expands the scope of the discussion to representations on Banach (as opposed to Hilbert) spaces $E$ in the sense of [8, Definition 3.1]: coassociative maps $E\xrightarrow{\rho}{\mathcal{C}}({\mathbb{G}})\otimes_{\lambda}E$ into the injective [12, §1.1] Banach-space tensor product satisfying the customary density condition (see ˜1-6 below).
 In that Banach-space context, said norm (or uniform) continuity is defined via [8, Lemma 3.7] as
@@
 One effect this will have is to render the $u^{\alpha}_{ij}$, (via [34, (5.14) and (5.15)] or [23, Theorem 1.4.3]), mutually orthogonal with respect to both versions of the inner product induced by the Haar state $h:=h_{{\mathbb{G}}}$ of [34, §4] or [23, §1.2]:
-(1-1) $\Braket{x\mid y}_{r}:=h(x^{*}y)\quad\text{and}\quad\Braket{x\mid y}_{\ell}:=h(xy^{*}),$
+(1-2) $\Braket{x\mid y}_{r}:=h(x^{*}y)\quad\text{and}\quad\Braket{x\mid y}_{\ell}:=h(xy^{*}),$
 the index indicating the position (left/right) of the product’s linear argument.
@@
 the $\sim$” indicating heuristic rather than an equality in any formal sense (convergence issues being ignored); the rigorous version is rather
-(1-2) $x^{\alpha}_{ij}:=\frac{q_{\alpha}}{F^{-1}_{\alpha ii}}(h\otimes\operatorname{id})\left(u^{\alpha*}_{ij}x\right)=\frac{q_{\alpha}}{F_{\alpha jj}}\left(h\otimes\operatorname{id})(xu^{\alpha*}_{ij}\right),$
+(1-3) $x^{\alpha}_{ij}:=\frac{q_{\alpha}}{F^{-1}_{\alpha ii}}(h\otimes\operatorname{id})\left(u^{\alpha*}_{ij}x\right)=\frac{q_{\alpha}}{F_{\alpha jj}}\left(h\otimes\operatorname{id})(xu^{\alpha*}_{ij}\right),$
 the $u^{\alpha}_{ij}$ being assumed $h$-orthogonal as previously announced. This suggests the equally informal
 $xx^{*}$ $\sim\sum_{\begin{subarray}{c}\alpha,i,j\\ \beta,k,l\end{subarray}}u^{\alpha}_{ij}u^{\beta*}_{kl}x^{\alpha}_{ij}x^{\beta*}_{kl}\xRightarrow{\quad}$
-(1-3) $E\left(xx^{*}\right):=\left(h\otimes\operatorname{id}\right)(xx^{*})$ $=\sum_{\begin{subarray}{c}\alpha,i,j\\ \beta,k,l\end{subarray}}h\left(u^{\alpha}_{ij}u^{\beta*}_{kl}\right)x^{\alpha}_{ij}x^{\beta*}_{kl}$
+(1-4) $E\left(xx^{*}\right):=\left(h\otimes\operatorname{id}\right)(xx^{*})$ $=\sum_{\begin{subarray}{c}\alpha,i,j\\ \beta,k,l\end{subarray}}h\left(u^{\alpha}_{ij}u^{\beta*}_{kl}\right)x^{\alpha}_{ij}x^{\beta*}_{kl}$
 […]
@@
 It remains to verify ˜c $\Rightarrow$ ˜a, which task the proof henceforth turns to.
-Reprising the notation of ˜1-3 for the usual conditional expectation [2, Definition II.6.10.1] ${\mathcal{C}}({\mathbb{G}})\underline{\otimes}{\mathcal{L}}(H)\xrightarrow[]{E}\mathrel{\mkern-14.0mu}\rightarrow{\mathcal{L}}(H)$, Proposition˜1.2 below ensures that that equation in fact holds as written, with a norm-convergent right-hand side.
-Applied to $x:=U$, ˜1-3 reads
+Reprising the notation of ˜1-4 for the usual conditional expectation [2, Definition II.6.10.1] ${\mathcal{C}}({\mathbb{G}})\underline{\otimes}{\mathcal{L}}(H)\xrightarrow[]{E}\mathrel{\mkern-14.0mu}\rightarrow{\mathcal{L}}(H)$, Proposition˜1.2 below ensures that that equation in fact holds as written, with a norm-convergent right-hand side.
+Applied to $x:=U$, ˜1-4 reads
 $1=E(1)=E\left(xx^{*}\right)$ $=\sum_{\alpha,j}\frac{1}{q_{\alpha}}F_{\alpha jj}\sum_{i}x^{\alpha}_{ij}x^{\alpha*}_{ij}$
@@
 as norm convergence, where
-(1-4) $P^{\alpha}:=(\psi^{\alpha}\otimes\operatorname{id})U^{*}\in{\mathcal{L}}(H),\quad{\mathcal{C}}({\mathbb{G}})\ni u^{\beta}_{ij}\xmapsto{\quad\psi^{\alpha}\quad}\delta_{\alpha\beta}\delta_{ii}\in{\mathbb{C}}$
+(1-5) $P^{\alpha}:=(\psi^{\alpha}\otimes\operatorname{id})U^{*}\in{\mathcal{L}}(H),\quad{\mathcal{C}}({\mathbb{G}})\ni u^{\beta}_{ij}\xmapsto{\quad\psi^{\alpha}\quad}\delta_{\alpha\beta}\delta_{ii}\in{\mathbb{C}}$
 are the respective projections (introduced in [25, Theorem 1.5] as $E^{\alpha}$) onto the $\alpha$-isotypic components (cf.
@@
   • $A\xrightarrow{\pi}{\mathcal{L}}(H)$ for a representation of $A$;
-  • ${\mathcal{C}}({\mathbb{G}})\xrightarrow{\pi_{h}}{\mathcal{L}}\left(L^{2}({\mathbb{G}})\right)$ for the GNS representation of the Haar state $h\in{\mathcal{C}}({\mathbb{G}})^{*}$: the completion of ${\mathcal{O}}({\mathbb{G}})$ under the inner product $\Braket{-\mid-}_{r}$ of ˜1-1;
+  • ${\mathcal{C}}({\mathbb{G}})\xrightarrow{\pi_{h}}{\mathcal{L}}\left(L^{2}({\mathbb{G}})\right)$ for the GNS representation of the Haar state $h\in{\mathcal{C}}({\mathbb{G}})^{*}$: the completion of ${\mathcal{O}}({\mathbb{G}})$ under the inner product $\Braket{-\mid-}_{r}$ of ˜1-2;
   • and
@@
 […]
-One again has spectral idempotents $P^{\alpha}:=(\psi^{\alpha}\otimes\operatorname{id})\circ\rho\in{\mathcal{L}}(E)$ analogous to ˜1-4 and, by [8, Proposition], the more familiar unitary representations $U\in M({\mathcal{C}}({\mathbb{G}})\underline{\otimes}{\mathcal{K}}(H))$ can be recast as representations on the underlying Banach spaces by
-(1-6) $H\ni v\xmapsto{\quad}U^{*}(1\otimes v)\in{\mathcal{C}}({\mathbb{G}})\otimes_{\lambda}H.$
+One again has spectral idempotents $P^{\alpha}:=(\psi^{\alpha}\otimes\operatorname{id})\circ\rho\in{\mathcal{L}}(E)$ analogous to ˜1-5 and, by [8, Proposition], the more familiar unitary representations $U\in M({\mathcal{C}}({\mathbb{G}})\underline{\otimes}{\mathcal{K}}(H))$ can be recast as representations on the underlying Banach spaces by
+(1-7) $H\ni v\xmapsto{\quad}U^{*}(1\otimes v)\in{\mathcal{C}}({\mathbb{G}})\otimes_{\lambda}H.$
 Proof of Theorem˜0.2.
@@
 Definition 1.3.
-A Banach-space representation ˜1-5 of a compact quantum group is boundedly uniform or uniform≤1 if the continuity in ˜0-1 holds on norm-bounded subsets of ${\mathcal{C}}({\mathbb{G}})^{*}$ (equivalently, on any one norm-bounded origin neighborhood, e.g.
+A Banach-space representation ˜1-6 of a compact quantum group is boundedly uniform or uniform≤1 if the continuity in ˜0-1 holds on norm-bounded subsets of ${\mathcal{C}}({\mathbb{G}})^{*}$ (equivalently, on any one norm-bounded origin neighborhood, e.g.
 the unit ball).
@@
 Proof.
-For the latter map this is immediate given the conversion procedure ˜1-6 from unitary to Banach-space representations, while the passage between the two is effected by $\left(\left(\psi\otimes\operatorname{id}\right)U\right)^{*}=\left(\psi^{*}\otimes\operatorname{id}\right)U^{*}$ for $\psi^{*}:=\overline{\psi\left(\bullet^{*}\right)}$.
+For the latter map this is immediate given the conversion procedure ˜1-7 from unitary to Banach-space representations, while the passage between the two is effected by $\left(\left(\psi\otimes\operatorname{id}\right)U\right)^{*}=\left(\psi^{*}\otimes\operatorname{id}\right)U^{*}$ for $\psi^{*}:=\overline{\psi\left(\bullet^{*}\right)}$.
 $\blacksquare$
@@
 The aforementioned symmetrization afforded by Proposition˜1.6 refers to switching Lemma˜1.4’s criterion to $U$’s right-hand leg. For $v\in H$ we write
-(1-7) $\operatorname{\mathrm{supp}}v:=\left\{\alpha\in\operatorname{Irr}({\mathbb{G}})\ :\ P_{U}^{\alpha}v\neq 0\right\}$
+(1-8) $\operatorname{\mathrm{supp}}v:=\left\{\alpha\in\operatorname{Irr}({\mathbb{G}})\ :\ P_{U}^{\alpha}v\neq 0\right\}$
 for the support of $v$ with respect to the representation $U$, and $\operatorname{\mathrm{supp}}v\to\infty$ for varying $v$ means that the support eventually misses every finite subset of $\operatorname{Irr}({\mathbb{G}})$.
@@
   1.
-˜1 Assume infinitely many spectral projections $P^{\alpha}$ of $U$ non-zero, entailing the corresponding non-vanishing Fourier coefficients $x^{\alpha}_{ij}$, $1\leq i,j\leq\dim\alpha$ of $x:=U$ defined by ˜1-2 (having fixed the $u^{\alpha}$-engendering bases of the $\alpha$ as explained preceding ˜1-1, so as to ensure the $u^{\alpha}_{ij}$ are mutually orthogonal).
+˜1 Assume infinitely many spectral projections $P^{\alpha}$ of $U$ non-zero, entailing the corresponding non-vanishing Fourier coefficients $x^{\alpha}_{ij}$, $1\leq i,j\leq\dim\alpha$ of $x:=U$ defined by ˜1-3 (having fixed the $u^{\alpha}$-engendering bases of the $\alpha$ as explained preceding ˜1-2, so as to ensure the $u^{\alpha}_{ij}$ are mutually orthogonal).
 The hypothesis ensures the existence of norm-1
@@
 as an algebra.
-  • The word length function [31, Example 3.2(ii)] $\ell$ on $\Gamma$ above with respect to a finite generating set ${\mathcal{F}}\subseteq\operatorname{Irr}({\mathbb{G}})$ is
+  • The word length function [31, Example 3.2(ii)] $\ell$ on $\Gamma$ above with respect to a finite generating set ${\mathcal{F}}\subseteq\operatorname{Irr}({\mathbb{G}})$ (typically assumed symmetric, i.e.
+with the dual $\alpha^{*}\in{\mathcal{F}}$ whenever $\alpha\in{\mathcal{F}}$) is
     $\operatorname{Irr}({\mathbb{G}})\ni\alpha\xmapsto{\quad\ell=\ell_{{\mathcal{F}}}\quad}\min\left\{k\in{\mathbb{Z}}_{\geq 0}\ :\ \exists\left(\alpha_{i}\in{\mathcal{F}},\ 1\leq i\leq k\right)\left(\alpha\leq\alpha_{1}\otimes\cdots\otimes\alpha_{k}\right)\right\}.$
@@
     for some polynomial $P$.
-  • The terminology applies here (somewhat unconventionally) also “locally” or relatively, to families of representations: given any finite ${\mathcal{F}}\subseteq\operatorname{Irr}({\mathbb{G}})$ one might consider sequences $\alpha_{n}$ exhibiting rapid decay with respect to $\ell:=\ell_{{\mathcal{F}}}$ in a sense the reader will have no difficulties extrapolating from the global version just sketched.
+  • The terminology applies here (somewhat unconventionally) also “locally” or relatively, to families of representations: given any symmetric finite ${\mathcal{F}}\subseteq\operatorname{Irr}({\mathbb{G}})$ one might consider sequences $\alpha_{n}$ exhibiting rapid decay with respect to $\ell:=\ell_{{\mathcal{F}}}$ in a sense the reader will have no difficulties extrapolating from the global version just sketched.
 The following result shows (e.g. via Example˜1.10) that there do exist uniform≤1 representations with infinite spectrum.
 Theorem 1.9.
-Let $\mathbb{G}$} be a compact quantum group, ${\mathcal{F}}\subseteq\operatorname{Irr}({\mathbb{G}})$ a finite family and $\left(\alpha_{n}\right)_{n}\subset\operatorname{Irr}({\mathbb{G}})$ with property RD with respect to $\ell_{{\mathcal{F}}}$ and trivial $F^{\alpha_{n}}$.
+Let $\mathbb{G}$} be a compact quantum group, ${\mathcal{F}}\subseteq\operatorname{Irr}({\mathbb{G}})$ a finite family and $\left(\alpha_{n}\right)_{n}\subset\operatorname{Irr}({\mathbb{G}})$ with property RD with respect to $\ell_{{\mathcal{F}}}$.
 If $\dim\alpha_{n}$ increase exponentially while the lengths $\ell\left(\alpha_{n}\right)$ increase polynomially, the unitary representation $\bigoplus_{n}\alpha_{n}$ is uniform≤1.
 Proof.
-Write $H:=\bigoplus_{n}H^{\alpha_{n}}$ for the carrier Hilbert space of the direct-sum representation $U$ in question. We have to argue, per Corollary˜1.7, that for $v\in H$ with
-(1-8) $\|v\|=1\ \wedge\ \operatorname{\mathrm{supp}}v\subseteq\left\{\alpha_{m}\ :\ m\gg 0\right\}$
-($\operatorname{\mathrm{supp}}$ defined as in ˜1-7) the norm $\left\|(\operatorname{id}\otimes\omega_{v,v})U\right\|$ will be as small as desired provided only that $m\gg 0$ is indeed sufficiently large.
-The unimodularity constraint $F^{\alpha}=1$ ensures [34, (5.14), (5.15)] that the $u^{\alpha_{n}}_{ij}$ (are mutually orthogonal and) have $\|-\|_{2}$ norms $\frac{1}{\sqrt{\dim\alpha_{n}}}$, hence the norm estimate
-˜1-8 $\xRightarrow{\quad}\ \exists\left(c>1\right)\left(\left\|(\operatorname{id}\otimes\omega_{v,v})U\right\|_{2}\leq\max_{m:\alpha_{m}\in\operatorname{\mathrm{supp}}v}\frac{1}{\sqrt{\dim\alpha_{m}}}\leq\frac{1}{c^{m_{0}}}\right)$
+Observe first that the individual positive operators $F_{\alpha_{n}}$ of ˜1-1 are identities: [31, Proposition 4.7] proves this globally, assuming the quantum group as a whole has property RD, but the argument (relying on [31, Lemmas 4.3 and 4.6]) goes through in the local version needed here.
+Write $H:=\bigoplus_{n}V_{\alpha_{n}}$ for the carrier Hilbert space of the direct-sum representation $U$ in question. We have to argue, per Corollary˜1.7, that for $v\in H$ with
+(1-9) $\|v\|=1\ \wedge\ \operatorname{\mathrm{supp}}v\subseteq\left\{\alpha_{m}\ :\ m\gg 0\right\}$
+($\operatorname{\mathrm{supp}}$ defined as in ˜1-8) the norm $\left\|(\operatorname{id}\otimes\omega_{v,v})U\right\|$ will be as small as desired provided only that $m\gg 0$ is indeed sufficiently large.
+The unimodularity condition $F_{\alpha}=1$ ensures [34, (5.14), (5.15)] that the $u^{\alpha_{n}}_{ij}$ (are mutually orthogonal and) have $\|-\|_{2}$ norms $\frac{1}{\sqrt{\dim\alpha_{n}}}$, hence the norm estimate
+˜1-9 $\xRightarrow{\quad}\ \exists\left(c>1\right)\left(\left\|(\operatorname{id}\otimes\omega_{v,v})U\right\|_{2}\leq\max_{m:\alpha_{m}\in\operatorname{\mathrm{supp}}v}\frac{1}{\sqrt{\dim\alpha_{m}}}\leq\frac{1}{c^{m_{0}}}\right)$
 $\xRightarrow{\quad\text{RD}\quad}\exists\left(\text{polynomial }P\right)\left(\left\|(\operatorname{id}\otimes\omega_{v,v})U\right\|\leq\frac{P(m_{0})}{c^{m_{0}}}\right)$

### 20 Mar 2026 v3

@@
   • on the other, representation-theoretic: the representation’s having finite spectrum, i.e. finitely many isotypic components.
-Classically, the two properties (norm continuity and spectral finiteness) are indeed equivalent: [15, p.257] addresses this for connected second-countable groups, [27, Corollary 2] in the broader setup of arbitrary compact groups represented norm-continuously on Banach spaces, and [8, Theorem 3.10] provides a number of equivalent characterizations.
+Classically, the two properties (norm continuity and spectral finiteness) are indeed equivalent: [15, p.257] addresses this for connected second-countable groups, [28, Corollary 2] in the broader setup of arbitrary compact groups represented norm-continuously on Banach spaces, and [9, Theorem 3.10] provides a number of equivalent characterizations.
 The compact quantum groups $\mathbb{G}$} to which the above uniformity/spectral finiteness discussion is to be extended are here as in [23, Definition 1.1.1]: unital $C^{*}$-algebras ${\mathcal{C}}({\mathbb{G}})$ (thought of as continuous-function algebras on the otherwise-nonexistent $\mathbb{G}$}) equipped with a coassociative unital $C^{*}$ morphism
@@
 There are alternative possible notions of norm continuity for compact-quantum-group representations, beyond those featuring in Theorem˜0.1.
-[8, Definition 3.6(3)], for instance, expands the scope of the discussion to representations on Banach (as opposed to Hilbert) spaces $E$ in the sense of [8, Definition 3.1]: coassociative maps $E\xrightarrow{\rho}{\mathcal{C}}({\mathbb{G}})\otimes_{\lambda}E$ into the injective [12, §1.1] Banach-space tensor product satisfying the customary density condition (see ˜1-6 below).
-In that Banach-space context, said norm (or uniform) continuity is defined via [8, Lemma 3.7] as
+[9, Definition 3.6(3)], for instance, expands the scope of the discussion to representations on Banach (as opposed to Hilbert) spaces $E$ in the sense of [9, Definition 3.1]: coassociative maps $E\xrightarrow{\rho}{\mathcal{C}}({\mathbb{G}})\otimes_{\varepsilon}E$ into the injective [27, §3.1] Banach-space tensor product satisfying the customary density condition (see ˜1-6 below).
+In that Banach-space context, said norm (or uniform) continuity is defined via [9, Lemma 3.7] as
 (0-1) ${\mathcal{C}}({\mathbb{G}})^{*}\ni\psi\xmapsto[\quad\text{weak${}^{*}$-to-norm continuous}\quad]{\quad}\left(v\xmapsto{\quad}(\psi\otimes\operatorname{id})\rho v\right)\in{\mathcal{L}}(E).$
@@
 Theorem 0.2.
-For a compact quantum group $\mathbb{G}$} represented via $E\xrightarrow{\rho}{\mathcal{C}}({\mathbb{G}})\otimes_{\lambda}E$ on a Banach space $E$ the following are equivalent.
+For a compact quantum group $\mathbb{G}$} represented via $E\xrightarrow{\rho}{\mathcal{C}}({\mathbb{G}})\otimes_{\varepsilon}E$ on a Banach space $E$ the following are equivalent.
   1. (a)
@@
     $\rho$ is norm-continuous in the sense that ˜0-1 holds.
-This reverses the implication recorded as [8, Lemma 3.8], but note that the hypothesis is arguably somewhat excessive: classically, one need not require the full power of ˜0-1, for in that case norm uniformity translates directly to the map in question being continuous when restricted to the space of pure states of ${\mathcal{C}}({\mathbb{G}})$ (i.e.
+This reverses the implication recorded as [9, Lemma 3.8], but note that the hypothesis is arguably somewhat excessive: classically, one need not require the full power of ˜0-1, for in that case norm uniformity translates directly to the map in question being continuous when restricted to the space of pure states of ${\mathcal{C}}({\mathbb{G}})$ (i.e.
 $\mathbb{G}$} itself, with points identified with their respective Dirac-delta measures).
@@
   1. (1)
-    If the Pontryagin dual [29, Theorem 3.3.2] $\Gamma:=\widehat{{\mathbb{G}}}$ has tempered decay in the sense that
-    $\begin{gathered}\exists\left(C>0\right)\forall\left(\alpha\in\operatorname{Irr}({\mathbb{G}})\right)\left(\left\|u^{\alpha}\right\|_{\infty}^{{\mathcal{C}}({\mathbb{G}})}>C\right)\\ \left\|u^{\alpha}\right\|_{\infty}^{{\mathcal{C}}({\mathbb{G}})}:=\sup_{\begin{subarray}{c}v,w\in{\mathbb{C}}^{\dim\alpha}\\ \|v\|,\|w\|\leq 1\end{subarray}}\left\|\Braket{v\mid u^{\alpha}\mid w}:=\sum_{i,j=1}^{\dim\alpha}u^{\alpha}_{ij}\overline{v_{i}}w_{j}\right\|.\end{gathered}$
-    $U$ is uniform≤1 if and only if it has finite spectrum.
+    If the Pontryagin dual [30, Theorem 3.3.2] $\Gamma:=\widehat{{\mathbb{G}}}$ has tempered decay in the sense that
+    $\begin{gathered}\exists\left(C>0\right)\forall\left(\alpha\in\operatorname{Irr}({\mathbb{G}})\right)\left(\left\|u^{\alpha}\right\|_{\infty}^{{\mathcal{C}}_{r}({\mathbb{G}})}>C\right)\\ \left\|u^{\alpha}\right\|_{\infty}^{{\mathcal{C}}({\mathbb{G}})}:=\sup_{\begin{subarray}{c}v,w\in{\mathbb{C}}^{\dim\alpha}\\ \|v\|,\|w\|\leq 1\end{subarray}}\left\|\Braket{v\mid u^{\alpha}\mid w}:=\sum_{i,j=1}^{\dim\alpha}u^{\alpha}_{ij}\overline{v_{i}}w_{j}\right\|,\end{gathered}$
+    then $U$ is uniform≤1 if and only if it has finite spectrum.
   2. (2)
-    In particular, said equivalence holds if $\mathbb{G}$} has a classical point, in the sense that ${\mathcal{C}}({\mathbb{G}})$ has a multiplicative state.
+    In particular, said equivalence holds if the reduced version of $\mathbb{G}$} has a classical point, in the sense that ${\mathcal{C}}_{r}({\mathbb{G}})$ has a multiplicative state.
+Equivalently: if $\mathbb{G}$} is coamenable [1, Definition 3.1].
 Counterexamples (see Theorem˜1.9) show that one cannot completely dispense with some type of decay control on the matrix coefficients of $\mathbb{G}$}.
 Acknowledgments
-Initial drafts have benefited from valuable input from M. Brannan.
+Initial drafts have benefited from M. Brannan’s valuable input
 1 Quantum uniform continuity in several guises, and consequences
-Some of the usual operator-algebraic notation and conventions will be operative: $A\bullet B$ with $\bullet\in\{\otimes,\ \underline{\otimes},\ \overline$$\otimes$}\} stand for the ordinary algebraic, and spatial (or minimal or injective) [28, Definitions IV.4.8 and IV.5.1] $C^{*}$ and $W^{*}$ tensor products respectively (between operator algebras of context-appropriate type), M()M(-) is the multiplier algebra [2, Theorem II.7.3.1], ${\mathcal{L}}(-,-)$ and ${\mathcal{K}}(-,-)$ denote spaces of bounded and compact operators respectively between Hilbert/Banach spaces, and we refer freely to slice maps ([2, §§II.9.7.1, III.2.2.6], [19, Lemma 1.5], [30, §1], etc.)
+Some of the usual operator-algebraic notation and conventions will be operative: $A\bullet B$ with $\bullet\in\{\otimes,\ \underline{\otimes},\ \overline$$\otimes$}\} stand for the ordinary algebraic, and spatial (or minimal or injective) [29, Definitions IV.4.8 and IV.5.1] $C^{*}$ and $W^{*}$ tensor products respectively (between operator algebras of context-appropriate type), M()M(-) is the multiplier algebra [3, Theorem II.7.3.1], ${\mathcal{L}}(-,-)$ and ${\mathcal{K}}(-,-)$ denote spaces of bounded and compact operators respectively between Hilbert/Banach spaces, and we refer freely to slice maps ([3, §§II.9.7.1, III.2.2.6], [19, Lemma 1.5], [31, §1], etc.)
 $\begin{aligned} M\left(A\underline{\otimes}B\right)&\xrightarrow{\quad\psi\otimes\operatorname{id}\quad}M(B)\\ A\overline{\otimes}B&\xrightarrow{\quad\psi\otimes\operatorname{id}\quad}B\end{aligned},\quad\psi\in A^{*}\left(:=\text{continuous dual of $A$}\right)$
-For compact quantum $\mathbb{G}$} with underlying function algebra ${\mathcal{C}}({\mathbb{G}})$ we work throughout with Woronowicz’s matrix coefficients ([34, §§1 and 4], [25, §1, pp.2-3], [23, §§1.3 and 1.4], etc.)
+For compact quantum $\mathbb{G}$} with underlying function algebra ${\mathcal{C}}({\mathbb{G}})$ we work throughout with Woronowicz’s matrix coefficients ([35, §§1 and 4], [25, §1, pp.2-3], [23, §§1.3 and 1.4], etc.)
 $u^{\alpha}_{ij}\in{\mathcal{O}}({\mathbb{G}}),\quad\begin{aligned} \alpha\in\operatorname{Irr}({\mathbb{G}})&:=\left\{\text{iso classes of simple ${\mathcal{O}}({\mathbb{G}})$-comodules}\right\}\\ 1\leq i,j\leq d_{\alpha}&:=\dim\alpha\end{aligned}$
 spanning the (unique dense [18, Theorem 3.1.7]) Hopf *-algebra of representative functions on a compact quantum group $\mathbb{G}$}.
-Unitary $\mathbb{G}$}-representations on Hilbert spaces $H$ are as in [11, §2.2]:
+Unitary $\mathbb{G}$}-representations on Hilbert spaces $H$ are as in [12, §2.2]:
 $\text{unitary }U\in M({\mathcal{C}}({\mathbb{G}})\underline{\otimes}{\mathcal{K}}(H)),\quad(\Delta\otimes\operatorname{id})U=U_{13}U_{23},$
@@
 diagonal.
-One effect this will have is to render the $u^{\alpha}_{ij}$, (via [34, (5.14) and (5.15)] or [23, Theorem 1.4.3]), mutually orthogonal with respect to both versions of the inner product induced by the Haar state $h:=h_{{\mathbb{G}}}$ of [34, §4] or [23, §1.2]:
+One effect this will have is to render the $u^{\alpha}_{ij}$, (via [35, (5.14) and (5.15)] or [23, Theorem 1.4.3]), mutually orthogonal with respect to both versions of the inner product induced by the Haar state $h:=h_{{\mathbb{G}}}$ of [35, §4] or [23, §1.2]:
 (1-2) $\Braket{x\mid y}_{r}:=h(x^{*}y)\quad\text{and}\quad\Braket{x\mid y}_{\ell}:=h(xy^{*}),$
 the index indicating the position (left/right) of the product’s linear argument.
+${\mathcal{C}}_{r}({\mathbb{G}})$ denotes the reduced version of $\mathcal{G}$}, i.e.
+[23, post Corollary 1.7.5] the image of ${\mathcal{C}}({\mathbb{G}})$ through the GNS representation attached to $h$.
 For a unital $C^{*}$-algebra $A$ an element $x\in{\mathcal{C}}({\mathbb{G}})\underline{\otimes}A$ has a Fourier-series decomposition
@@
 […]
-The extent to which this last equality holds formally, with the right-hand side norm-convergent, is the focus of amply studied Fourier decay-at-infinity phenomena: quantum analogues of the familiar [4, §I.2] Riemann-Lebesgue lemma in classical Fourier analysis, as alluded to briefly in [1, §1, p.34], say.
+The extent to which this last equality holds formally, with the right-hand side norm-convergent, is the focus of amply studied Fourier decay-at-infinity phenomena: quantum analogues of the familiar [5, §I.2] Riemann-Lebesgue lemma in classical Fourier analysis, as alluded to briefly in [2, §1, p.34], say.
 We revisit the matter in the course of the proof of Theorem˜0.1.
@@
 Proof of Theorem˜0.1.
-The equivalence ˜a $\Leftrightarrow$ ˜b is covered by [8, Theorem 3.16], while ˜a is plainly stronger than the other conditions: finite spectrum in fact entails
+The equivalence ˜a $\Leftrightarrow$ ˜b is covered by [9, Theorem 3.16], while ˜a is plainly stronger than the other conditions: finite spectrum in fact entails
 $U\in{\mathcal{O}}({\mathbb{G}})\otimes{\mathcal{L}}(H)\quad\left(\text{algebraic tensor product}\right).$
 It remains to verify ˜c $\Rightarrow$ ˜a, which task the proof henceforth turns to.
-Reprising the notation of ˜1-4 for the usual conditional expectation [2, Definition II.6.10.1] ${\mathcal{C}}({\mathbb{G}})\underline{\otimes}{\mathcal{L}}(H)\xrightarrow[]{E}\mathrel{\mkern-14.0mu}\rightarrow{\mathcal{L}}(H)$, Proposition˜1.2 below ensures that that equation in fact holds as written, with a norm-convergent right-hand side.
+Reprising the notation of ˜1-4 for the usual conditional expectation [3, Definition II.6.10.1] ${\mathcal{C}}({\mathbb{G}})\underline{\otimes}{\mathcal{L}}(H)\xrightarrow[]{E}\mathrel{\mkern-14.0mu}\rightarrow{\mathcal{L}}(H)$, Proposition˜1.2 below ensures that that equation in fact holds as written, with a norm-convergent right-hand side.
 Applied to $x:=U$, ˜1-4 reads
@@
 are the respective projections (introduced in [25, Theorem 1.5] as $E^{\alpha}$) onto the $\alpha$-isotypic components (cf.
-[3, remarks preceding Proposition 13]) of the representation $U$.
+[4, remarks preceding Proposition 13]) of the representation $U$.
 Said norm convergence of mutually orthogonal projections, naturally, requires that only finitely many be non-zero.
 $\blacksquare$
-The proof of the auxiliary Proposition˜1.2 below will precisely parallel its analogue [26, Lemma 5.2] in the context of reduced crossed products [6, Definition 4.1.4] resulting from actions by countable (classical, discrete) groups on $C^{*}$-algebras (so in our language, that result’s $\mathbb{G}$} would be an abelian, i.e.
+The proof of the auxiliary Proposition˜1.2 below will precisely parallel its analogue [26, Lemma 5.2] in the context of reduced crossed products [7, Definition 4.1.4] resulting from actions by countable (classical, discrete) groups on $C^{*}$-algebras (so in our language, that result’s $\mathbb{G}$} would be an abelian, i.e.
 dual-classical, compact quantum group).
@@
 $\sum_{\alpha,i,j}\frac{q_{\alpha}}{F_{\alpha jj}}T^{\alpha}_{ij}T^{\alpha*}_{ij}=1\in{\mathcal{L}}\left(L^{2}({\mathbb{G}})\otimes H\right)$
-in the $\sigma$-strong topology of [28, Definition II.2.3] (the strongest of the locally convex sub-norm topologies examined in [28, §II.2]): the individual summands $\frac{q_{\alpha}}{F_{\alpha jj}}T^{\alpha}_{ij}T^{\alpha*}_{ij}$ are (by [34, (5.14)] again) mutually orthogonal projections whose ranges exhaust the underlying Hilbert space.
+in the $\sigma$-strong topology of [29, Definition II.2.3] (the strongest of the locally convex sub-norm topologies examined in [29, §II.2]): the individual summands $\frac{q_{\alpha}}{F_{\alpha jj}}T^{\alpha}_{ij}T^{\alpha*}_{ij}$ are (by [35, (5.14)] again) mutually orthogonal projections whose ranges exhaust the underlying Hilbert space.
 Note also:
@@
 […]
-One again has spectral idempotents $P^{\alpha}:=(\psi^{\alpha}\otimes\operatorname{id})\circ\rho\in{\mathcal{L}}(E)$ analogous to ˜1-5 and, by [8, Proposition], the more familiar unitary representations $U\in M({\mathcal{C}}({\mathbb{G}})\underline{\otimes}{\mathcal{K}}(H))$ can be recast as representations on the underlying Banach spaces by
-(1-7) $H\ni v\xmapsto{\quad}U^{*}(1\otimes v)\in{\mathcal{C}}({\mathbb{G}})\otimes_{\lambda}H.$
+One again has spectral idempotents $P^{\alpha}:=(\psi^{\alpha}\otimes\operatorname{id})\circ\rho\in{\mathcal{L}}(E)$ analogous to ˜1-5 and, by [9, Proposition], the more familiar unitary representations $U\in M({\mathcal{C}}({\mathbb{G}})\underline{\otimes}{\mathcal{K}}(H))$ can be recast as representations on the underlying Banach spaces by
+(1-7) $H\ni v\xmapsto{\quad}U^{*}(1\otimes v)\in{\mathcal{C}}({\mathbb{G}})\otimes_{\varepsilon}H.$
 Proof of Theorem˜0.2.
-While [8, Lemma 3.8] notes only that having finite spectrum implies ˜0-1, the converse is also straightforward: weak-to-norm continuity is an extremely strong condition to impose, holding only in trivial circumstances.
+While [9, Lemma 3.8] notes only that having finite spectrum implies ˜0-1, the converse is also straightforward: weak-to-norm continuity is an extremely strong condition to impose, holding only in trivial circumstances.
 To elaborate, consider a linear map $E\xrightarrow{T}F$ between locally convex [16, §18.1] topological vector spaces. Its continuity with respect to
@@
 $\blacklozenge$
-For classical $\mathbb{G}$} bounded and plain uniformity are equivalent: as noted in the Introduction following Theorem˜0.2, uniformity≤1 already implies spectrum finiteness and hence [8, Lemma 3.8] also ordinary uniformity.
+For classical $\mathbb{G}$} bounded and plain uniformity are equivalent: as noted in the Introduction following Theorem˜0.2, uniformity≤1 already implies spectrum finiteness and hence [9, Lemma 3.8] also ordinary uniformity.
 It is not inapposite, then, to ask whether the equivalence survives quantization.
-As further motivation and surrounding context for Theorem˜0.3, recall [7, Remark 1.3] to the effect that the classical [27, Corollary 2] characterization of norm-continuous compact-group representations as precisely those with finite spectra is recovered by non-classical means from [7, Theorem 0.1] for compact quantum groups $\mathbb{G}$} with a classical point (i.e.
+As further motivation and surrounding context for Theorem˜0.3, recall [8, Remark 1.3] to the effect that the classical [28, Corollary 2] characterization of norm-continuous compact-group representations as precisely those with finite spectra is recovered by non-classical means from [8, Theorem 0.1] for compact quantum groups $\mathbb{G}$} with a classical point (i.e.
 with ${\mathcal{C}}({\mathbb{G}})$ admitting a multiplicative state); by extension, the same goes for Theorem˜0.1 (which removes the classical-point constraint).
@@
 $\blacksquare$
-The device isolated by Proposition˜1.6 below, implicitly also driving the proof of [30, Proposition 2.2], will allow some symmetrization in the condition on $U$ imposed by Lemma˜1.4.
+The device isolated by Proposition˜1.6 below, implicitly also driving the proof of [31, Proposition 2.2], will allow some symmetrization in the condition on $U$ imposed by Lemma˜1.4.
 Some vocabulary will help compress the statement.
@@
   4. (d)
-    As in ˜a, for nets of functionals of the form
-    $M\ni T\xmapsto{\quad\omega_{v,w}\quad}\Braket{v\mid Tw}\in{\mathbb{C}}$
-    for $v,w\in H$ in a Hilbert space on which $M$ is represented faithfully as $M\leq{\mathcal{L}}(H)$.
+    Setting $\omega_{v,w}:=\Braket{v\mid\bullet w}\in M_{*}$ for $v,w\in H$ in a Hilbert space on which $M$ is represented faithfully as $M\leq{\mathcal{L}}(H)$,
+    $\omega_{v,w}\xmapsto{\quad}(\omega_{v,w}\otimes\operatorname{id})x\in{\mathcal{C}}({\mathbb{G}})$
+    turns weak-Cauchy nets into their norm counterparts.
   5. (e)
@@
 Note that ˜a concerns all of $M^{*}$, whereas the other $\bullet_{\ell}$ specialize to (at least) weak-continuous functionals.
-The proof of [30, Proposition 2.2], in present language, shows effectively that the strongest left-handed statement ˜a is implied by its right-handed counterpart phrased for nets $(\psi_{\lambda})_{\lambda}\subset N_{*,\leq 1}$ (unit ball of the predual of $N$) assumed convergent $N^{*}_{\leq 1}$.
-This is easily recast as ˜br though: the net in question is certainly Cauchy and hence $(\varphi_{\lambda}-\varphi_{\lambda^{\prime}})_{\lambda,\lambda^{\prime}}$ is 0-convergent; consequently,
+The proof of [31, Proposition 2.2], in present language, shows effectively that the strongest left-handed statement ˜a is implied by its right-handed counterpart phrased for nets $(\psi_{\lambda})_{\lambda}\subset N_{*,\leq 1}$ (unit ball of the predual of $N$) assumed convergent $N^{*}_{\leq 1}$.
+This is easily recast as ˜br though: the net in question is certainly Cauchy and hence $(\psi_{\lambda}-\psi_{\lambda^{\prime}})_{\lambda,\lambda^{\prime}}$ is 0-convergent; consequently,
 […]
-˜c $\Rightarrow$ ˜b by weak-in-weak norm density [28, Theorem II.2.6(iii)].
-In order to complete the proof of the proposition as a whole by confirming that ˜er implies ˜a, it remains to observe that the aforementioned net $(\psi_{\lambda})_{\lambda}\subset N_{*,\leq 1}$ can be chosen so as to range over any set ${\mathcal{S}}\subset N_{\leq 1}^{*}$ detecting norms in $N$ in the following sense:
-$\forall\left(T\in N\right)\forall\left(\varepsilon>0\right)\exists\left(\varphi\in{\mathcal{S}}\right)\left(|\varphi T|\geq\|T\|-\varepsilon\right).$
+˜c $\Rightarrow$ ˜b by weak-in-weak norm density [29, Theorem II.2.6(iii)].
+In order to complete the proof of the proposition as a whole by confirming that ˜dr implies ˜a, it remains to observe that the aforementioned net $(\psi_{\lambda})_{\lambda}\subset N_{*,\leq 1}$ can be chosen so as to range over any set ${\mathcal{S}}\subset N_{\leq 1}^{*}$ detecting norms in $N$ in the following sense:
+$\forall\left(T\in N\right)\forall\left(\varepsilon>0\right)\exists\left(\psi\in{\mathcal{S}}\right)\left(|\psi T|\geq\|T\|-\varepsilon\right).$
 Plainly, the set $\mathcal{S}$} of all functionals $\omega_{v,w}$ with norm-$(\leq 1)$ $v,w\in H$ will do. $\blacksquare$
-The aforementioned symmetrization afforded by Proposition˜1.6 refers to switching Lemma˜1.4’s criterion to $U$’s right-hand leg. For $v\in H$ we write
-(1-8) $\operatorname{\mathrm{supp}}v:=\left\{\alpha\in\operatorname{Irr}({\mathbb{G}})\ :\ P_{U}^{\alpha}v\neq 0\right\}$
-for the support of $v$ with respect to the representation $U$, and $\operatorname{\mathrm{supp}}v\to\infty$ for varying $v$ means that the support eventually misses every finite subset of $\operatorname{Irr}({\mathbb{G}})$.
-Corollary 1.7.
-A unitary representation $U\in M({\mathcal{C}}({\mathbb{G}})\underline{\otimes}{\mathcal{K}}(H))$ of a compact quantum group $\mathbb{G}$} is uniform≤1 as a Banach-space representation if and only if
-$\left(\operatorname{\mathrm{supp}}v\to\infty\wedge\operatorname{\mathrm{supp}}w\to\infty\right)\ \xRightarrow{\quad}\ (\operatorname{id}\otimes\omega_{v,w})U\xrightarrow[\quad\quad]{\quad\text{norm}\quad}0$
-for $(v,w)\in\left(H_{\leq 1}\right)^{2}$ and $\omega_{v,w}:=\Braket{v\mid\bullet w}$.
-Proof.
-Lemma˜1.4 and Proposition˜1.6’s equivalence ˜a $\Leftrightarrow$ ˜dr equate uniformity≤1 with the weak-norm convergence of
-$\omega_{v,w}\xmapsto{\quad}\left(\operatorname{id}\otimes\omega_{v,w}\right)U\in{\mathcal{L}}(H),\quad v,w\in H_{\leq 1}.$
-We may thus assume weakly convergent nets
-$\left(v_{\lambda}\right)_{\lambda}\xrightarrow[\quad\lambda\quad]{\quad}v\quad\text{and}\quad\left(w_{\lambda}\right)_{\lambda}\xrightarrow[\quad\lambda\quad]{\quad}w$
-in the unit ball $H_{\leq 1}$, whence $\omega_{v_{\lambda},w_{\lambda}}\to\omega_{v,w}$.
-Approximate $v$ and $w$ arbitrarily well by finitely supported vectors $\overline{v}$ and $\overline{w}$ respectively, assuming also that
-$\operatorname{\mathrm{supp}}\left(v_{\lambda}-\overline{v}\right)\to\infty\quad\text{and}\quad\operatorname{\mathrm{supp}}\left(w_{\lambda}-\overline{w}\right)\to\infty.$
-Because finite truncations
-$\sum_{\alpha\in{\mathcal{F}}}\left(1\otimes P^{\alpha}\right)U,\quad\text{finite }{\mathcal{F}}\subseteq\operatorname{Irr}({\mathbb{G}})$
-belong to the algebraic tensor product ${\mathcal{O}}({\mathbb{G}})\otimes{\mathcal{L}}(H)$, the norm convergence of the truncated
-$\left(\operatorname{id}\otimes\omega_{\bullet}\right)U,\quad\bullet\in\left\{(\overline{v},\overline{w}),\ \left(v_{\lambda}-\overline{v},\overline{w}\right),\ \left(\overline{v},w_{\lambda}-\overline{w}\right)\right\}$
-is not at issue. This confirms the equivalence
-$\left(\operatorname{id}\otimes\omega_{v_{\lambda},w_{\lambda}}\right)U\xrightarrow[\quad\lambda\quad]{\quad\text{norm}\quad}\left(\operatorname{id}\otimes\omega_{v,w}\right)U\quad\xLeftrightarrow{\quad}\quad\left(\operatorname{id}\otimes\omega_{v_{\lambda}-\overline{v},w_{\lambda}-\overline{w}}\right)U\xrightarrow[\quad\lambda\quad]{\quad\text{norm}\quad}0,$
-finishing the proof. $\blacksquare$
+The aforementioned symmetrization afforded by Proposition˜1.6 refers to switching Lemma˜1.4’s criterion to $U$’s right-hand leg, as will become apparent presently.
 Proof of Theorem˜0.3.
@@
 ˜1 Assume infinitely many spectral projections $P^{\alpha}$ of $U$ non-zero, entailing the corresponding non-vanishing Fourier coefficients $x^{\alpha}_{ij}$, $1\leq i,j\leq\dim\alpha$ of $x:=U$ defined by ˜1-3 (having fixed the $u^{\alpha}$-engendering bases of the $\alpha$ as explained preceding ˜1-2, so as to ensure the $u^{\alpha}_{ij}$ are mutually orthogonal).
-The hypothesis ensures the existence of norm-1
-    $v_{0}=\left(v_{0i}\right)_{1\leq i\leq\dim\alpha}\quad\text{and}\quad w_{0}=\left(w_{0i}\right)_{1\leq i\leq\dim\alpha}\quad:\quad\left\|\sum_{i,j=1}^{\dim\alpha}u^{\alpha}_{ij}\overline{v_{0i}}w_{0j}\right\|>C.$
-    To procure suitable $\omega_{v,w}$ amenable to Corollary˜1.7, set
-    $v:=\sum_{i=1}^{\dim\alpha}v_{0i}e_{i}\quad\text{and}\quad w:=\sum_{i=1}^{\dim\alpha}w_{0i}e_{i},\quad e_{i}\in\operatorname{\mathrm{im}}x^{\alpha}_{ii}\leq\operatorname{\mathrm{im}}P^{\alpha}\text{ orthonormal}.$
-    We have
-    $\left\|(\operatorname{id}\otimes\omega_{v,w})U\right\|=\left\|\sum_{i,j=1}^{\dim\alpha}u^{\alpha}_{ij}\overline{v_{0i}}w_{0j}\right\|>C,$
-    affording the desired conclusion that uniformity≤1 must fail.
+The tempered-decay hypothesis ensures the existence of norm-1
+    $v_{\alpha 0}=\left(v_{\alpha 0i}\right)_{1\leq i\leq\dim\alpha}\quad\text{and}\quad w_{\alpha 0}=\left(w_{\alpha 0i}\right)_{1\leq i\leq\dim\alpha}\quad:\quad\left\|\sum_{i,j=1}^{\dim\alpha}u^{\alpha}_{ij}\overline{v_{\alpha 0i}}w_{\alpha 0j}\right\|>C.$
+    Now consider $\omega_{\alpha}:=\omega_{v_{\alpha},w_{\alpha}}:=\Braket{v_{\alpha}\mid\bullet w_{\alpha}}\in{\mathcal{L}}(H)_{*}$ for
+    $v_{\alpha}:=\sum_{i=1}^{\dim\alpha}v_{\alpha 0i}e_{i}\quad\text{and}\quad w_{\alpha}:=\sum_{i=1}^{\dim\alpha}w_{\alpha 0i}e_{i},\quad e_{i}\in\operatorname{\mathrm{im}}x^{\alpha}_{ii}\leq\operatorname{\mathrm{im}}P^{\alpha}\text{ orthonormal},$
+    so that by construction we have
+    $\left\|(\operatorname{id}\otimes\omega_{\alpha})U\right\|=\left\|\sum_{i,j=1}^{\dim\alpha}u^{\alpha}_{ij}\overline{v_{\alpha 0i}}w_{\alpha 0j}\right\|>C.$
+    By Alaoglu’s [10, Theorem V.3.1], the bounded family $\{\omega_{\alpha}\}_{\alpha}$ supports a weak-Cauchy net which we index by the $\alpha$ themselves, with the understanding that $\alpha$ eventually avoids every finite subset of $\operatorname{Irr}({\mathbb{G}})$.
+    Now, on the one hand the net $\left((\operatorname{id}\otimes\omega_{\alpha})U\right)_{\alpha}$ must norm-converge in ${\mathcal{C}}_{r}({\mathbb{G}})$ to an element of norm $\geq C>0$.
+On the other, matrix-coefficient Haar orthogonality forces
+[…]
+    $h$ being faithful [23, Corollary 1.7.5] on ${\mathcal{C}}_{r}({\mathbb{G}})$, the desired conclusion that uniformity≤1 must fail follows from that property’s characterization as Proposition˜1.6’s right-handed ˜d.
   2.
-˜2 $\mathbb{G}$} has a classical point precisely when the counit $\varepsilon$ has a continuous extension to all of ${\mathcal{C}}({\mathbb{G}})$ (see [10, paragraph preceding Theorem 2.2]; also Lemma˜1.8 below for additional equivalent characterizations), whence a state assigning the value 11 to every $u^{\alpha}_{ii}$.
-Those elements, then, must all have norm 1.
+˜2 $\mathbb{G}$} has a classical point precisely when the counit $\varepsilon$ has a continuous extension to all of ${\mathcal{C}}({\mathbb{G}})$ (see [11, paragraph preceding Theorem 2.2]; also Lemma˜1.7 below for additional equivalent characterizations), whence a state on ${\mathcal{C}}_{r}({\mathbb{G}})$ assigning the value $1$ to every $u^{\alpha}_{ii}$.
+Those elements, then, must all have norm 1 and the conclusion follows from part ˜1.
 $\blacksquare$
-In reference to the property featuring in [7, Theorem 0.1] of compact quantum groups having at least one classical point (one version of the above-mentioned “partial classical behavior”), Lemma˜1.8 records a simple remark for whatever future-reference purpose it may serve.
+In reference to the property featuring in [8, Theorem 0.1] of compact quantum groups having at least one classical point (one version of the above-mentioned “partial classical behavior”), Lemma˜1.7 records a simple remark for whatever future-reference purpose it may serve.
 We will be referring to the usual convolution product [21, Definition 1.4.1], denoted here simply by $\cdot$” or juxtaposition, and defined as the precomposition
@@
 on the algebraic dual of a coalgebra.
-Note also that it (or rather its appropriate extension) makes the state space ${\mathcal{S}}({\mathcal{C}}({\mathbb{G}}))$ of the $C^{*}$-algebra ${\mathcal{C}}({\mathbb{G}})$ into a compact Hausdorff [9, Theorem V.3.1] semigroup under its weak topology.
-The same goes for the unit ball ${\mathcal{C}}({\mathbb{G}})^{*}_{\leq 1}$ in the continuous dual of ${\mathcal{C}}({\mathbb{G}})$: the norm dual to $\underline{\otimes}$ being cross [32, Proposition T.3.11],
+Note also that it (or rather its appropriate extension) makes the state space ${\mathcal{S}}({\mathcal{C}}({\mathbb{G}}))$ of the $C^{*}$-algebra ${\mathcal{C}}({\mathbb{G}})$ into a compact Hausdorff [10, Theorem V.3.1] semigroup under its weak topology.
+The same goes for the unit ball ${\mathcal{C}}({\mathbb{G}})^{*}_{\leq 1}$ in the continuous dual of ${\mathcal{C}}({\mathbb{G}})$: the norm dual to $\underline{\otimes}$ being cross [33, Proposition T.3.11],
 $\|\varphi\|,\ \|\psi\|\leq 1\xRightarrow{\quad}\|\varphi\cdot\psi:=(\varphi\otimes\psi)\circ\Delta\|\leq 1.$
@@
 That the variants of the claim are mutually equivalent follows from the fact that arbitrary continuous functionals are expressible as linear combinations of at most four states (e.g.
-by [28, Theorem III.4.2(ii)]).
+by [29, Theorem III.4.2(ii)]).
 Having thus reduced the problem to its ${\mathcal{C}}({\mathbb{G}})^{*}$ variant, write $\psi$ for the assumed unit of that convolution monoid and
@@
     The $\varepsilon_{{\mathcal{F}}}$ being restrictions to ${\mathcal{O}}({\mathbb{G}})$ of continuous functionals (i.e.
-members of ${\mathcal{C}}({\mathbb{G}})^{*}$) by [34, (5.37) and surrounding discussion],
+members of ${\mathcal{C}}({\mathbb{G}})^{*}$) by [35, (5.37) and surrounding discussion],
     $\psi=\lim_{{\mathcal{F}}\nearrow}\psi\cdot\varepsilon_{{\mathcal{F}}}=\lim_{{\mathcal{F}}\nearrow}\varepsilon_{{\mathcal{F}}}=\varepsilon$
@@
     $\psi^{{n_{\lambda,\lambda^{\prime}}}}\cdot\in\left(\textsc{Cont}\left({\mathcal{K}}\to{\mathcal{K}}\right),\ \text{uniform topology}\right)$
-    for $\mathcal{K}$}: this follows from the identification [5, Proposition 7.1.5]
+    for $\mathcal{K}$}: this follows from the identification [6, Proposition 7.1.5]
     $\textsc{Cont}\left({\mathcal{K}}\times{\mathcal{K}}\to{\mathcal{K}}\right)\cong\textsc{Cont}\left({\mathcal{K}}\to\textsc{Cont}\left({\mathcal{K}}\to{\mathcal{K}}\right)\right).$
@@
 as an algebra.
-  • The word length function [31, Example 3.2(ii)] $\ell$ on $\Gamma$ above with respect to a finite generating set ${\mathcal{F}}\subseteq\operatorname{Irr}({\mathbb{G}})$ (typically assumed symmetric, i.e.
+  • The word length function [32, Example 3.2(ii)] $\ell$ on $\Gamma$ above with respect to a finite generating set ${\mathcal{F}}\subseteq\operatorname{Irr}({\mathbb{G}})$ (typically assumed symmetric, i.e.
 with the dual $\alpha^{*}\in{\mathcal{F}}$ whenever $\alpha\in{\mathcal{F}}$) is
@@
     We extend the notation $\ell(-)$ to elements of ${\mathcal{O}}({\mathbb{G}})$: one such has length $k$ by definition if it is a linear combination of matrix coefficients $u^{\alpha}_{ij}$ with $\ell(\alpha)=k$.
-  • $\Gamma$ has the property of rapid decay (RD) [31, Proposition and Definition 3.5] (with respect to $\ell_{{\mathcal{F}}}$) if
+  • $\Gamma$ has the property of rapid decay (RD) [32, Proposition and Definition 3.5] (with respect to $\ell_{{\mathcal{F}}}$) if
     $\left(\forall x\in{\mathcal{O}}({\mathbb{G}})\right)\left(\|x\|\leq P(\ell(x))\|x\|_{2}\right),\quad\|-\|_{2}:=h\left(-^{*}\cdot-\right)^{1/2}$
@@
   • The terminology applies here (somewhat unconventionally) also “locally” or relatively, to families of representations: given any symmetric finite ${\mathcal{F}}\subseteq\operatorname{Irr}({\mathbb{G}})$ one might consider sequences $\alpha_{n}$ exhibiting rapid decay with respect to $\ell:=\ell_{{\mathcal{F}}}$ in a sense the reader will have no difficulties extrapolating from the global version just sketched.
-The following result shows (e.g. via Example˜1.10) that there do exist uniform≤1 representations with infinite spectrum.
+Theorem˜1.9 shows (e.g. via Example˜1.10) that there do exist uniform≤1 representations with infinite spectrum; an auxiliary remark precedes it.
+For $v\in H$ we write
+(1-8) $\operatorname{\mathrm{supp}}v:=\left\{\alpha\in\operatorname{Irr}({\mathbb{G}})\ :\ P_{U}^{\alpha}v\neq 0\right\}$
+for the support of $v$ with respect to the representation $U$, and $\operatorname{\mathrm{supp}}v\to\infty$ for varying $v$ means that the support eventually misses every finite subset of $\operatorname{Irr}({\mathbb{G}})$.
+Lemma 1.8.
+A unitary representation $U\in M({\mathcal{C}}({\mathbb{G}})\underline{\otimes}{\mathcal{K}}(H))$ of a compact quantum group $\mathbb{G}$} is uniform≤1 as a Banach-space representation provided
+$\left(\operatorname{\mathrm{supp}}v\to\infty\wedge\operatorname{\mathrm{supp}}w\to\infty\right)\ \xRightarrow{\quad}\ (\operatorname{id}\otimes\omega_{v,w})U\xrightarrow[\quad\quad]{\quad\text{norm}\quad}0$
+for $(v,w)\in\left(H_{\leq 1}\right)^{2}$ and $\omega_{v,w}:=\Braket{v\mid\bullet w}$.
+Proof.
+Per Lemma˜1.4, we have to verify Proposition˜1.6’s condition ˜a.
+A slightly more careful reworking of the implication ˜dr $\Rightarrow$ ˜a (building on the proof of [31, Proposition 2.2]) will deliver that outcome given the present hypotheses.
+Recall to that end that said proof begins by assuming a weak-0-convergent net $\left(\varphi_{\lambda}\right)_{\lambda}\subset{\mathcal{C}}({\mathbb{G}})^{*}$ with $U_{\varphi_{\lambda}}:=\left(\varphi_{\lambda}\otimes\operatorname{id}\right)U$ norm-bounded away from 0 and selecting corresponding
+$\psi_{\lambda}:=\omega_{v_{\lambda},w_{\lambda}}\in{\mathcal{L}}(H)_{*},\quad v_{\lambda},w_{\lambda}\in H_{\leq 1}$
+with $\left|\psi_{\lambda}U_{\varphi_{\lambda}}\right|$ sufficiently close to $\left\|U_{\varphi_{\lambda}}\right\|$. Write
+$P^{{\mathcal{F}}}:=\sum_{\alpha\in{\mathcal{F}}}P^{\alpha},\quad{\mathcal{F}}\subseteq\operatorname{Irr}({\mathbb{G}})$
+The hypothesis ensures that for finite ${\mathcal{F}}\subseteq\operatorname{Irr}({\mathbb{G}})$ sufficiently large,
+$\left|\psi^{{\mathcal{F}}}_{\lambda}U_{\varphi_{\lambda}}-\psi_{\lambda}U_{\varphi_{\lambda}}\right|$ $=\left|\varphi_{\lambda}\left(\operatorname{id}\otimes\psi^{{\mathcal{F}}}_{\lambda}\right)U-\varphi_{\lambda}\left(\operatorname{id}\otimes\psi_{\lambda}\right)U\right|$
+$\psi^{{\mathcal{F}}}_{\lambda}$ $\in\left\{\omega_{P^{{\mathcal{F}}}v_{\lambda},w_{\lambda}},\ \omega_{v_{\lambda},P^{{\mathcal{F}}}w_{\lambda}},\ \omega_{P^{{\mathcal{F}}}v_{\lambda},P^{{\mathcal{F}}}w_{\lambda}}\right\}$
+is $\lambda$-uniformly small, so we may assume all $v_{\lambda}$ and $w_{\lambda}$ finitely supported on a single ${\mathcal{F}}\subseteq\operatorname{Irr}({\mathbb{G}})$.
+The truncation $\left(1\otimes P^{{\mathcal{F}}}\right)U$ belonging to the algebraic tensor product
+${\mathcal{C}}({\mathbb{G}})^{{\mathcal{F}}}\otimes{\mathcal{L}}(H),\quad{\mathcal{C}}({\mathbb{G}}):=\bigoplus_{\begin{subarray}{c}\alpha\in{\mathcal{F}}\\ 1\leq i,j\leq d_{\alpha}\end{subarray}}{\mathbb{C}}u^{\alpha}_{ij},$
+the weak-norm Cauchy preservation assumed in the proof of Proposition˜1.6’s ˜dr $\Rightarrow$ ˜a implication holds for our $\psi_{\lambda}$ and the argument goes through. $\blacksquare$
 Theorem 1.9.
@@
 Proof.
-Observe first that the individual positive operators $F_{\alpha_{n}}$ of ˜1-1 are identities: [31, Proposition 4.7] proves this globally, assuming the quantum group as a whole has property RD, but the argument (relying on [31, Lemmas 4.3 and 4.6]) goes through in the local version needed here.
-Write $H:=\bigoplus_{n}V_{\alpha_{n}}$ for the carrier Hilbert space of the direct-sum representation $U$ in question. We have to argue, per Corollary˜1.7, that for $v\in H$ with
+We will apply Lemma˜1.8, proving that its hypothesis obtains under the present circumstances.
+Observe first that the individual positive operators $F_{\alpha_{n}}$ of ˜1-1 are identities: [32, Proposition 4.7] proves this globally, assuming the quantum group as a whole has property RD, but the argument (relying on [32, Lemmas 4.3 and 4.6]) goes through in the local version needed here.
+Writing $H:=\bigoplus_{n}V_{\alpha_{n}}$ for the carrier Hilbert space of the direct-sum representation $U$ in question, we have to argue that for $v\in H$ with
 (1-9) $\|v\|=1\ \wedge\ \operatorname{\mathrm{supp}}v\subseteq\left\{\alpha_{m}\ :\ m\gg 0\right\}$
 ($\operatorname{\mathrm{supp}}$ defined as in ˜1-8) the norm $\left\|(\operatorname{id}\otimes\omega_{v,v})U\right\|$ will be as small as desired provided only that $m\gg 0$ is indeed sufficiently large.
-The unimodularity condition $F_{\alpha}=1$ ensures [34, (5.14), (5.15)] that the $u^{\alpha_{n}}_{ij}$ (are mutually orthogonal and) have $\|-\|_{2}$ norms $\frac{1}{\sqrt{\dim\alpha_{n}}}$, hence the norm estimate
+The unimodularity condition $F_{\alpha}=1$ ensures [35, (5.14), (5.15)] that the $u^{\alpha_{n}}_{ij}$ (are mutually orthogonal and) have $\|-\|_{2}$ norms $\frac{1}{\sqrt{\dim\alpha_{n}}}$, hence the norm estimate
 ˜1-9 $\xRightarrow{\quad}\ \exists\left(c>1\right)\left(\left\|(\operatorname{id}\otimes\omega_{v,v})U\right\|_{2}\leq\max_{m:\alpha_{m}\in\operatorname{\mathrm{supp}}v}\frac{1}{\sqrt{\dim\alpha_{m}}}\leq\frac{1}{c^{m_{0}}}\right)$
@@
 Examples of compact quantum groups verifying the hypotheses of Theorem˜1.9 are well known and have been much studied (from various perspectives, RD-related or not): both
-  • the free orthogonal group ${\mathbb{G}}:={\mathbb{O}}^{+}(n)$, $n\geq 3$ (underlying function algebra ${\mathcal{C}}({\mathbb{G}})=A_{o}(I_{n})$ in [31, §4.3]), with ${\mathcal{O}}({\mathbb{G}})$ freely generated by a unitary $n\times n$ matrix $u^{\alpha}=\left(u^{\alpha}_{ij}\right)$, $1\leq i,j\leq n$ of self-adjoint elements;
+  • the free orthogonal group ${\mathbb{G}}:={\mathbb{O}}^{+}(n)$, $n\geq 3$ (underlying function algebra ${\mathcal{C}}({\mathbb{G}})=A_{o}(I_{n})$ in [32, §4.3]), with ${\mathcal{O}}({\mathbb{G}})$ freely generated by a unitary $n\times n$ matrix $u^{\alpha}=\left(u^{\alpha}_{ij}\right)$, $1\leq i,j\leq n$ of self-adjoint elements;
   • and the free unitary group ${\mathbb{G}}:={\mathbb{U}}^{+}(n)$, $n\geq 3$, with ${\mathcal{O}}({\mathbb{G}})$ freely generated by unitary $u^{\alpha}\in{\mathcal{O}}({\mathbb{G}})\otimes M_{n}$ such that $\overline{u}^{\alpha}:=\left(u^{\alpha*}_{ij}\right)_{i,j}$ is also unitary
-will do by [31, Theorems 4.9 and 4.10] respectively. $\blacklozenge$
+will do by [32, Theorems 4.9 and 4.10] respectively. $\blacklozenge$